### v1

# On Fock covariance for product systems and the reduced Hao–Ng isomorphism problem by discrete actions

Abstract.
We provide a characterisation of equivariant Fock covariant injective representations for product systems. We show that this characterisation coincides with Nica covariance for compactly aligned product systems over right LCM semigroups of Kwaśniewski and Larsen, and with the Toeplitz representations of a discrete monoid of Laca and Sehnem. By combining with the framework established by Katsoulis and Ramsey, we resolve the reduced Hao–Ng isomorphism problem for generalised gauge actions by discrete groups.
2020 Mathematics Subject Classification. 46L08, 47L55, 46L05
Key words and phrases: Product systems, Fock space, crossed products of operator algebras.
1. Introduction
1.1. Fock covariance
Hilbertian representations play an important role in the study of algebraic structures and symmetries. One of the cornerstone examples is the Gel’fand–Raikov Theorem that “identifies” a locally compact group with its unitary representations. Further examples come from C*-dynamical systems, (topological) graphs, semigroups, and groupoids. Algebraic structures associated with discrete groups have been successfully unified under the theory of Fell bundles. Product systems over a unital subsemigroup $P$ of a discrete group $G$ provide a common context to describe irreversible transformations as a semigroup analogue of Fell bundles. There has been a great number of results in the area, perhaps too many to list here, and we will just comment on work as related to our results.
Product systems were introduced by Arveson [2] disguised under duality, and later put in context by Dinh [10] for discrete subsemigroups of $\mathbb{R}_{+}$. Following Nica [42] and Pimsner [45], Fowler [18] proceeded to an in-depth study when $(G,P)$ is a quasi-lattice, that inspired a great number of subsequent works. It had been known that, unlike to the group case, taking plain isometric representations leads to intractable objects, e.g., the universal isometric C*-algebra of $\mathbb{N}\times\mathbb{N}$ is not even nuclear [40]. A richer structure is at hand when one considers the relations in the Fock representation of a quasi-lattice order, known as Nica covariance. In order to make full use of this additional feature, Fowler imposed several axioms on the product system resulting to a by-default Wick ordering. Kwaśniewski and Larsen [32, 33] extended this setup to the case of right LCM semigroups, and essentially this is the furthest one may go and still have a Wick ordering. Cuntz, Deninger and Laca [8] used the Fock model beyond Nica’s work [42] for $P=R\rtimes R^{\times}$, where Nica covariance may fail, but one can still use the Fock representation as the prototype for covariant relations.
The coaction of the ambient group $G$ has been pivotal in the theory of product systems, in particular in relation to boundary quotients by Carslen, Larsen, Sims and Vittadello [7] and Sehnem [47, 48], as well as in semigroup algebras by Laca and Sehnem [34] and the first author, Katsoulis, Laca and Li [22]. Dor-On, the first author, Katsoulis, Laca and Li [11] noted that there is a canonical Fell bundle arising from the normal coaction on ${\mathcal{T}}_{\lambda}(X)$, denoted here by ${\mathcal{F}}{\mathcal{C}}X$. It appears that, while ${\mathcal{T}}_{\lambda}(X)$ is the reduced C*-algebra of ${\mathcal{F}}{\mathcal{C}}X$, it is not always the reduced C*-algebra for the universal C*-algebra ${\mathcal{T}}(X)$ of the representations of $X$ (as it happens for $P=\mathbb{Z}_{+}$). To make a distinction, we will write ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ for the universal C*-algebra of ${\mathcal{F}}{\mathcal{C}}X$; a representation of $X$ that promotes to a *-representation of ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ will be called Fock covariant. It was shown in [11] that Fock covariant representations coincide with Nica covariant representations for a compactly aligned product system $X$ over a right LCM semigroup $P$, and thus ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ coincides with the universal Nica covariant C*-algebra ${\mathcal{N}}{\mathcal{T}}(X)$ in this case.
The difference between ${\mathcal{T}}(X)$ and ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ is apparent even when every $X$ arises from $P$ by one-dimensional spaces, where the unital representations of $X$ are in bijection with the unital semigroup representations of $P$. In his ground-breaking work, Li [36] took motivation from [8] and revived the interest on semigroup representations with the twist that, apart from the semigroup structure, they remember the principal ideals and their intersections, i.e., the constructible ideals. Li’s models are variants of the relations in the Fock representation and were further examined by the first author, Katsoulis, Laca and Li [22], also in connection with inverse semigroup realizations by Norling [43]. At the same time Laca and Sehnem [34] identified completely the relations that promote a general semigroup representation to a Fock covariant representation of its related product system. It is thus natural to ask for the characterisation of Fock covariance for general product systems.
Motivated by the Hao–Ng isomorphism problem, our first aim in this work is to provide a description of the equivariant Fock covariant injective representations of $X$. For our purposes, we will view a product system $X$ as a family of C*-correspondences $\{X_{p}\}_{p\in P}$ in a common ${\mathcal{B}}(H)$ that satisfies some natural semigroup relations, i.e., that $A:=X_{e}$ is a C*-algebra, and that
(1.1) $X_{p}\cdot X_{q}\subseteq X_{pq}\text{ and }X_{p}^{*}\cdot X_{pq}\subseteq X_{q}\text{ for all }p,q\in P.$
This setup enables to still define left creation operators on the full Fock space ${\mathcal{F}}X$ giving rise to the Fock C*-algebra ${\mathcal{T}}_{\lambda}(X)$, while it recovers Fowler’s product systems in [18] which arise under the stronger saturation condition that $[X_{p}X_{q}]=X_{pq}$ for all $p,q\in P$. We note that most of the works for product systems assume non-degeneracy of the left action on $X$. Nevertheless, non-degeneracy is rarely used, although imposed by the saturation condition of Fowler [18]. The characterisation we provide is given in terms of the $\mathbf{K}_{{\mathbf{x}}}$-cores for ${\mathbf{x}}\in{\mathcal{J}}$. We then use the natural description of the fixed point algebra by an inductive limit of C*-algebras on the $\cap$-closed finite subsets of the constructible ideals of $P$ in ${\mathcal{J}}$, which are spanned by the $\mathbf{K}_{\bullet}$-cores.
Theorem A. (Theorem 3.2) Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\hat{t}$ be a representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. An equivariant injective representation $t$ of $X$ is Fock covariant if and only if $t$ satisfies the following conditions:
  1. (i)
    $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$.
  2. (ii)
    For any $\cap$-closed =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}\subseteq{\mathcal{J}}$ such that $\bigcup_{i=1}^{n}\mathbf{x}_{i}\neq\emptyset$, and any $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$, with $i=1,\dots,n$, the following property holds:
    if $\sum\limits_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(X_{r})=(0)$ for all $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then $\sum\limits_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=0$.
We square this characterisation with established results from the literature. Towards this end, we provide a direct connection with the equivariant Nica covariant injective representations when $P$ is a right LCM semigroup recovering the key result of [11], and with the equivariant Toeplitz semigroup representations of Laca and Sehnem [34] when $X$ arises from $P$ (note that injectivity is automatic in the latter case). As a consequence we get that Fock covariance coincides with Nica covariance for compactly aligned product systems over a right LCM semigroup, and with Toeplitz covariance for semigroup representations.
As we will see, a description of Fock covariance just for the case of equivariant representations that are injective on $X$ is enough to tackle key questions in the theory. One application of particular interest, and our main motivation for this work, concerns the reduced Hao–Ng isomorphism problem [20] which we describe below.
1.2. The Hao–Ng isomorphism problem
Along with ${\mathcal{T}}(X)$ and ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ there is a significant boundary quotient. For the case $P=\mathbb{Z}_{+}$ this is the Cuntz–Pimsner algebra ${\mathcal{O}}_{X}$, which generalises the graph C*-algebra and the Cuntz algebra, and was provided in full generality by Katsura [29] following the work of many hands. Katsura [30] has also proven that ${\mathcal{O}}_{X}$ is the terminal object for the equivariant injective representations of $X$ (Fock covariance is automatically satisfied in this case), while Katsoulis and Kribs [26] have shown that ${\mathcal{O}}_{X}$ is the C*-envelope of the tensor algebra of $X$. Carlsen, Larsen, Sims and Vittadello [7] have provided the existence of the terminal object in the case of a quasi-lattice, while Dor-On, the first author, Katsoulis, Laca and Li [11] have tackled the case for right LCM semigroups in terms of a terminal object for the coaction on the tensor algebra. A pivotal step in the general direction was established by Sehnem [47] where the appropriate universal quotient $A\times_{X}P$ was coined. The key property of the equivariant representations of $A\times_{X}P$ is that injectivity on $X$ implies injectivity on the fixed point algebra. Equivariance allows to view $A\times_{X}P$ as the universal C*-algebra of the strong covariant bundle denoted by ${\mathcal{S}}{\mathcal{C}}X$ in [11]. The terminal object in [11] is then the reduced C*-algebra of ${\mathcal{S}}{\mathcal{C}}X$ denoted by $A\times_{X,\lambda}P$. Sehnem [48] has shown that completely isometric representations of the tensor algebra ${\mathcal{T}}_{\lambda}(X)^{+}$ admit automatically a conditional expectation. With this remarkable result at hand, Sehnem [48] resolved the old standing problem of identifying $A\times_{X,\lambda}P$ with the C*-envelope of ${\mathcal{T}}_{\lambda}(X)^{+}$, encompassing the work of many authors.
The reduced Hao–Ng isomorphism problem refers to a group action $\alpha$ of ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$ that leaves invariant every $\lambda_{p}(X_{p})$; such an action is called a generalised gauge action. In this case there is an induced product system denoted by $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ and an induced action $\dot{\alpha}$ on $A\times_{X,\lambda}P$. The main question is whether there is a canonical *-isomorphism
(1.2) $(A\rtimes_{\alpha,\lambda}{\mathfrak{H}})\times_{X\rtimes_{\alpha,\lambda}{\mathfrak{H}},\lambda}P\stackrel{{\scriptstyle?}}{{\simeq}}(A\times_{X,\lambda}P)\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}},$
i.e., if the reduced strong covariant functor commutes with the reduced crossed product functor. The problem was introduced in [20], where it was solved for $P=\mathbb{Z}_{+}$ and ${\mathfrak{H}}$ being an amenable locally compact group. As applications of their results, Hao and Ng recover previous results on Hilbert bimodules by Abadie [1], and on graph C*-algebras by Kumjian and Pask [31] without using groupoid C*-algebras. The importance of the Hao–Ng isomorphism problem has been further emphasised by Katsoulis [24] in connection to the work of Echterhoff, Kaliszewski, Quigg and Raeburn [15] on imprimitivity theorems for C*-dynamical systems.
In a series of works, Katsoulis [24, 25], and Katsoulis and Ramsey [28], have rephrased the reduced Hao–Ng isomorphism problem in terms of C*-envelopes and crossed products [27], i.e., whether there exists a canonical *-isomorphism
(1.3) $\mathrm{C}^{*}_{\textup{env}}({\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{+})\stackrel{{\scriptstyle?}}{{\simeq}}\mathrm{C}^{*}_{\textup{env}}({\mathcal{T}}_{\lambda}(X)^{+})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}.$
The significance of this approach is manifested in [24] where the problem was resolved for $P=\mathbb{Z}_{+}$ and ${\mathfrak{H}}$ discrete. The reduced Hao–Ng isomorphism problem was later answered when $P$ defines an abelian lattice order in $G$ by Dor-On and Katsoulis [12] when ${\mathfrak{H}}$ is discrete, and by Katsoulis [25] when ${\mathfrak{H}}$ is locally compact abelian. Furthermore it has been answered when $P$ is a right LCM semigroup and ${\mathfrak{H}}$ is discrete by Dor-On, the first author, Katsoulis, Laca and Li [11]. The question remains open for a generalised gauge action by a general group ${\mathfrak{H}}$.
The main approach in [11, 12, 24, 25, 28] has been to use the independence condition for right LCM semigroups and Nica covariance of the identity representation
(1.4) $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\hookrightarrow{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$
in order to obtain a canonical *-isomorphism
(1.5) ${\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})\simeq{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
From there it follows that
(1.6) ${\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{+}\simeq{\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}},$
and then the C*-envelope theory of Katsoulis and Ramsey [27] can be implemented. In the absence of independence, it is unclear whether the *-isomorphism (1.5) still holds, but it does not exclude the possibility that the completely isometric isomorphism (1.6) is valid. A careful investigation of the arguments of [11, 12, 24, 25, 28] reveals that Fock covariance of the identity representation $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\hookrightarrow{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ enables the completely isometric isomorphism (1.6). Our characterisation applies towards this resolution when ${\mathfrak{H}}$ is discrete without restrictions on the product system $X$ and the semigroup $P$.
Theorem B. (Theorem 4.6) Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\alpha$ be a generalised gauge action of a discrete group ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$. Then the identity representation $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\hookrightarrow{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ lifts to a completely isometric isomorphism
${\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{+}\simeq{\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
Consequently, the reduced Hao–Ng isomorphism problem has an affirmative answer, i.e.,
$(A\rtimes_{\alpha,\lambda}{\mathfrak{H}})\times_{X\rtimes_{\alpha,\lambda}{\mathfrak{H}},\lambda}P\simeq(A\times_{X,\lambda}P)\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}},$
by a canonical *-isomorphism, where $\dot{\alpha}$ is the induced action of ${\mathfrak{H}}$ on $A\times_{X,\lambda}P$.
The structure of the manuscript is as follows. In Section 2 we fix notation and provide some relevant results we will be using. In Section 3 we give the description of Fock covariance, and applications to compactly aligned product systems over right LCM semigroups in connection to [11, 32, 33], as well as to semigroup representations in connection to [34]. In Section 4 we provide the context for the reduced Hao–Ng isomorphism problem, and then proceed to its resolution when ${\mathfrak{H}}$ is discrete. Throughout, we comment when non-degeneracy is not required, and we provide details when a different proof is needed, when using results from the literature.
Acknowledgements.
The authors would like to thank Elias Katsoulis for his helpful comments and remarks on the draft of the manuscript. The second author acknowledges support from CaLISTA which funded the research visit of the second author to Newcastle University in October 2023. This manuscript will form part of the second author’s PhD thesis.
Open Access Statement.
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript (AAM) version arising.
2. Preliminaries
We begin with some preliminaries on coactions on C*-algebras, semigroup C*-algebras and product systems to fix notation. All groups and semigroups we consider in this section are discrete. We will write $\otimes$ for the minimal tensor product between C*-algebras. A map between algebras will be called canonical if it preserves generators of the same index. For notational convenience, we will write $x^{0}=1$ for an element $x$ in a unital algebra with unit $1$. If $E$ is a subset of a normed linear space $F$ then we will write $[E]$ for the closed linear span generated by $E$ in $F$.
2.1. Coactions on C*-algebras
For a discrete group $G$ we write $u_{g}$ for the generators of the universal group C*-algebra $\mathrm{C}^{*}_{\max}(G)$, and $\lambda_{g}:=\lambda(u_{g})$ for the left regular representation $\lambda\colon\mathrm{C}^{*}_{\max}(G)\to\mathrm{C}^{*}_{\lambda}(G)$. We write $\chi$ for the character on $\mathrm{C}^{*}_{\max}(G)$. By the universal property of $\mathrm{C}^{*}_{\max}(G)$ there exists a faithful *-homomorphism
$\Delta\colon\mathrm{C}^{*}_{\max}(G)\to\mathrm{C}^{*}_{\max}(G)\otimes\mathrm{C}^{*}_{\max}(G);u_{g}\mapsto u_{g}\otimes u_{g}.$
By Fell’s absorption principle there exists a faithful *-homomorphism
$\Delta_{\lambda}\colon\mathrm{C}^{*}_{\lambda}(G)\to\mathrm{C}^{*}_{\lambda}(G)\otimes\mathrm{C}^{*}_{\lambda}(G);\lambda_{g}\mapsto\lambda_{g}\otimes\lambda_{g},$
with the additional property that $\Delta_{\lambda}\circ\lambda=(\lambda\otimes\lambda)\circ\Delta$.
There is a direct connection between coactions on C*-algebras, gradings and Fell bundles. The reader is addressed to [16, 17, 46] where this theory is laid in full detail. Since the group $G$ is discrete, the coactions we consider are automatically non-degenerate in the sense of [46], see for example [11, Remark 3.2].
We will say that a C*-algebra ${\mathcal{C}}$ admits a coaction $\delta$ by $G$ if there is a faithful *-homomorphism $\delta\colon{\mathcal{C}}\to{\mathcal{C}}\otimes\mathrm{C}^{*}_{\max}(G)$ such that the coaction identity
$(\delta\otimes{\operatorname{id}})\circ\delta=({\operatorname{id}}\otimes\Delta)\circ\delta$
is satisfied. Since $G$ is discrete, the coaction identity is equivalent to the induced spectral spaces
${\mathcal{C}}_{g}:=\{c\in{\mathcal{C}}\mid\delta(c)=c\otimes u_{g}\}\text{ forall }g\in G,$
together spanning a dense subset of ${\mathcal{C}}$, see the proof of [41, Proposition 2.6]. If, in addition, the map $({\operatorname{id}}\otimes\lambda)\circ\delta$ is faithful, then $\delta$ will be called normal. It follows that $\delta$ is normal if and only if ${\mathcal{C}}$ admits a reduced coaction by $G$, i.e., there is a faithful *-homomorphism $\delta_{\lambda}\colon{\mathcal{C}}\to{\mathcal{C}}\otimes\mathrm{C}^{*}_{\lambda}(G)$ such that
$(\delta_{\lambda}\otimes{\operatorname{id}})\circ\delta_{\lambda}=({\operatorname{id}}\otimes\Delta_{\lambda})\circ\delta_{\lambda}.$
We note that if $\delta\colon{\mathcal{C}}\to{\mathcal{C}}\otimes\mathrm{C}^{*}_{\max}(G)$ is a coaction, then
$E:=({\operatorname{id}}\otimes E_{\lambda})\circ({\operatorname{id}}\otimes\lambda)\circ\delta$
defines a conditional expectation on ${\mathcal{C}}_{e}$, where $E_{\lambda}$ is the faithful conditional expectation of $\mathrm{C}^{*}_{\lambda}(G)$. It follows that $\delta$ is normal if and only if $E$ is faithful.
More generally, a collection $\{{\mathcal{C}}_{g}\}_{g\in G}$ of closed subspaces of a C*-algebra ${\mathcal{C}}$ is called a C*-grading of ${\mathcal{C}}$ if:
  1. (i)
    $\sum_{g\in G}{\mathcal{C}}_{g}$ is dense in ${\mathcal{C}}$;
  2. (ii)
    ${\mathcal{C}}_{g}{\mathcal{C}}_{h}\subseteq{\mathcal{C}}_{gh}$; and
  3. (iii)
    ${\mathcal{C}}_{g}^{*}\subseteq{\mathcal{C}}_{g^{-1}}$.
If there exists a conditional expectation $E\colon{\mathcal{C}}\to{\mathcal{C}}_{e}$, then the subspaces are linearly independent, see [16, Theorem 3.3]. In this case the C*-grading is called topological. By definition, a coaction on a C*-algebra induces a topological C*-grading.
Gradings form the prototypical example of Fell bundles. A Fell bundle over a group $G$ is a collection =${\mathcal{B}}=\{{\mathcal{B}}_{g}\}_{g\in G}$ of Banach spaces, each of which is called a fiber, such that:
  1. (i)
    there are bilinear and associative multiplication maps from ${\mathcal{B}}_{g}\times{\mathcal{B}}_{g^{\prime}}$ to ${\mathcal{B}}_{gg^{\prime}}$ such that $\|b_{g}b_{g^{\prime}}\|\leq\|b_{g}\|\,\|b_{g^{\prime}}\|$;
  2. (ii)
    there are conjugate linear involution maps from ${\mathcal{B}}_{g}$ to ${\mathcal{B}}_{g^{-1}}$ such that $(b_{g}^{*})^{*}=b_{g}$ and $\|b_{g}^{*}\|=\|b_{g}\|$;
  3. (iii)
    $(b_{g}b_{g^{\prime}})^{*}=(b_{g^{\prime}})^{*}b_{g}^{*}$;
  4. (iv)
    $\|b_{g}^{*}b_{g}\|=\|b_{g}\|^{2}$;
  5. (v)
    $b_{g}^{*}b_{g}\geq 0$ in ${\mathcal{B}}_{e}$.
Note that conditions (i)–(iv) imply that ${\mathcal{B}}_{e}$ is in fact a C*-algebra and thus condition (v) makes sense. A representation $\Psi$ of a Fell bundle ${\mathcal{B}}$ over $G$ is a family $\{\Psi_{g}\}_{g\in G}$ of linear maps each one defined on ${\mathcal{B}}_{g}$ such that:
  1. (i)
    $\Psi_{g}(b_{g})\Psi_{h}(b_{h})=\Psi_{gh}(b_{g}b_{h})$ for all $g,h\in G$; and
  2. (ii)
    $\Psi_{g}(b_{g})^{*}=\Psi_{g^{-1}}(b_{g}^{*})$ for all $g\in G$.
It follows that $\Psi_{e}$ is a *-homomorphism and thus contractive. A standard C*-trick shows that every $\Psi_{g}$ is contractive. We say that a representation $\Psi$ is injective if $\Psi_{e}$ is injective; in this case every $\Psi_{g}$ is isometric. A representation $\Psi$ is called equivariant if there exists a *-homomorphism
$\delta\colon\mathrm{C}^{*}(\Psi)\to\mathrm{C}^{*}(\Psi)\otimes\mathrm{C}^{*}(G);\Psi_{g}(b_{g})\mapsto\Psi_{g}(b_{g})\otimes u_{g}.$
It follows that $\delta$ is faithful, with a left inverse given by the map ${\operatorname{id}}\otimes\chi$, and that it satisfies the coaction identity.
We write $\mathrm{C}^{*}_{\max}({\mathcal{B}})$ for the universal C*-algebra with respect to the representations of ${\mathcal{B}}$ and we write
$\hat{j}\colon{\mathcal{B}}\to\mathrm{C}^{*}_{\max}({\mathcal{B}})$
for the canonical embedding. We use the same symbol $\Psi$ for the *-homomorphism of $\mathrm{C}^{*}_{\max}({\mathcal{B}})$ induced by a representation $\{\Psi_{g}\}_{g\in G}$ of ${\mathcal{B}}$ (as $\Psi\circ\hat{j}_{g}=\Psi_{g}$). By universality, we have that $\{\hat{j}_{g}\}_{g\in G}$ is an equivariant representation of ${\mathcal{B}}$, and in particular by [17, Proposition 17.9] we have that the map $\hat{j}_{g}\colon{\mathcal{B}}_{g}\to[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}$ is an isometric isomorphism. Hence, any Fell bundle arises as a C*-grading from some C*-algebra.
The left regular representation of a Fell bundle ${\mathcal{B}}$ is defined by considering the left creation operators
$(\lambda_{g}(b_{g})\xi)_{g^{\prime}}=b_{g}\xi_{g^{-1}g^{\prime}}\text{ for all}b_{g}\in{\mathcal{B}},$
on $\ell^{2}({\mathcal{B}}):=\sum_{g\in G}^{\oplus}{\mathcal{B}}_{g}$ seen as the Hilbert module direct sum over ${\mathcal{B}}_{e}$. We write $\mathrm{C}^{*}_{\lambda}({\mathcal{B}})$ for the C*-algebra generated by $\lambda$.
By writing $j\colon{\mathcal{B}}\to\ell^{2}({\mathcal{B}})$ for the canonical embedding of each fiber at the corresponding summand of $\ell^{2}({\mathcal{B}})$, we can define the unitary
$W\colon\ell^{2}({\mathcal{B}})\otimes\ell^{2}(G)\to\ell^{2}({\mathcal{B}})\otimes\ell^{2}(G);j_{g}(b_{g})\otimes\delta_{g^{\prime}}\mapsto j_{g}(b_{g})\otimes\delta_{gg^{\prime}}.$
It follows that $W$ implements a reduced coaction
$\mathrm{C}^{*}_{\lambda}({\mathcal{B}})\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\mathrm{C}^{*}(\lambda_{g}(b_{g})\otimes I)\stackrel{{\scriptstyle\operatorname{ad}_{W}}}{{\longrightarrow}}\mathrm{C}^{*}_{\lambda}({\mathcal{B}})\otimes\mathrm{C}^{*}_{\lambda}(G),$
and thus a normal coaction. By [16, Theorem 3.3] we have that, if $\Psi$ is an equivariant representation of ${\mathcal{B}}$ that is injective on ${\mathcal{B}}_{e}$, then there are equivariant canonical *-epimorphisms
$\mathrm{C}^{*}_{\max}({\mathcal{B}})\longrightarrow\mathrm{C}^{*}(\Psi)\longrightarrow\mathrm{C}^{*}_{\lambda}({\mathcal{B}}).$
If, in addition, the coaction on $\mathrm{C}^{*}(\Psi)$ is normal, then $\mathrm{C}^{*}(\Psi)\simeq\mathrm{C}^{*}_{\lambda}({\mathcal{B}})$.
More generally, let $\Psi$ be an equivariant (possibly non-injective) representation of ${\mathcal{B}}$, then by the proof of [17, Proposition 21.4] (or by using the Fourier transform) we obtain that every $\Psi_{g}$ has closed range, and hence we have the induced Fell bundle
$\Psi({\mathcal{B}}):=\{\Psi_{g}({\mathcal{B}}_{g})\}_{g\in G}.$
Therefore there are equivariant *-epimorphisms, making the following diagram
${\mathrm{C}^{*}_{\max}({\mathcal{B}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{C}^{*}_{\lambda}({\mathcal{B}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{C}^{*}_{\max}(\Psi({\mathcal{B}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{C}^{*}(\Psi)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{C}^{*}_{\lambda}(\Psi({\mathcal{B}}))}$
commutative, see [17, Proposition 21.2, Proposition 21.3]. We will make use of the following folklore fact for Fell bundles.
Proposition 2.1.
Let ${\mathcal{B}}$ be a Fell bundle over a discrete group $G$, and $\Psi_{1}$ and $\Psi_{2}$ be equivariant representations of ${\mathcal{B}}$. Then
$\ker\Psi_{1}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e}\subseteq\ker\Psi_{2}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e}$
if and only if
$\ker\Psi_{1}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}\subseteq\ker\Psi_{2}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}\text{ for all }g\in G.$
If any of the above holds, and $\Psi_{1}({\mathcal{B}})$ is the induced Fell bundle in $\mathrm{C}^{*}(\Psi_{1})$, then $\Psi_{2}$ defines a representation of $\Psi_{1}({\mathcal{B}})$, and thus there is a commutative diagram
${\mathrm{C}^{*}_{\max}({\mathcal{B}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{C}^{*}(\Psi_{2})}$${\mathrm{C}^{*}_{\max}(\Psi_{1}({\mathcal{B}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
of equivariant *-epimorphisms.
Proof..
It is obvious that if the inclusion holds for all $g\in G$ then in particular it holds for $g=e$. Conversely, suppose that
$\ker\Psi_{1}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e}\subseteq\ker\Psi_{2}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e},$
and let $x\in\ker\Psi_{1}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}$ for $g\in G$. Then
$x^{*}x\in\ker\Psi_{1}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e}\subseteq\ker\Psi_{2}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e},$
and so $x\in\ker\Psi_{2}$. Consequently we derive $x\in\ker\Psi_{2}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{g}$, as required.
For the second part of the proof, the existence of the maps
$\mathrm{C}^{*}_{\max}({\mathcal{B}})\to\mathrm{C}^{*}_{\max}(\Psi_{1}({\mathcal{B}}))\quad\text{and}\quad\mathrm{C}^{*}_{\max}({\mathcal{B}})\to\mathrm{C}^{*}(\Psi_{2})$
follows by the discussion prior to the statement and the universal property of the universal C*-algebras. Due to the inclusion of the kernels, for every $g\in G$ we have a commutative diagram
${{\mathcal{B}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{2}|_{{\mathcal{B}}_{g}}}$$\scriptstyle{\Psi_{1}|_{{\mathcal{B}}_{g}}}$${\Psi_{2}({\mathcal{B}}_{g})}$${\Psi_{1}({\mathcal{B}}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{g}}$
such that $\Psi_{g}(\Psi_{1}(b_{g}))=\Psi_{2}(b_{g})$ for every $b_{g}\in{\mathcal{B}}_{g}$, and $g\in G$. Since $\Psi_{1}$ and $\Psi_{2}$ are representations of ${\mathcal{B}}$, we get that the collection $\{\Psi_{g}\}_{g\in G}$ defines a representation $\Psi$ from $\Psi_{1}({\mathcal{B}})$ to $\mathrm{C}^{*}(\Psi_{2})$. Hence $\Psi$ promotes to a *-representation of $\mathrm{C}^{*}_{\max}(\Psi_{1}({\mathcal{B}}))$ in $\mathrm{C}^{*}(\Psi_{2})$. By definition this map closes the diagram, and the proof is complete. ∎
We will be interested in graded quotients of C*-algebras of Fell bundles. If $\delta\colon{\mathcal{C}}\to{\mathcal{C}}\otimes\mathrm{C}^{*}_{\max}(G)$ is a coaction on a C*-algebra ${\mathcal{C}}$, then we say that an ideal ${\mathcal{I}}\lhd{\mathcal{C}}$ is induced if
${\mathcal{I}}=\left\langle{\mathcal{I}}\cap[{\mathcal{C}}]_{e}\right\rangle.$
In that case, the canonical quotient map $q_{\mathcal{I}}$ is equivariant, i.e., the coaction $\delta$ descends to a coaction on ${\mathcal{C}}/{\mathcal{I}}$, see [16, Proposition 3.11] and [7, Proposition A.1] for the full details. The following proposition is perhaps folklore, and we include a proof for completeness.
Proposition 2.2.
Let ${\mathcal{B}}$ be a Fell bundle over a discrete group $G$ and let ${\mathcal{I}}\lhd\mathrm{C}^{*}_{\max}({\mathcal{B}})$ be an induced ideal. Let $q_{\mathcal{I}}({\mathcal{B}})$ be the induced Fell bundle from the coaction on $\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}}$, where $q_{\mathcal{I}}\colon\mathrm{C}^{*}_{\max}({\mathcal{B}})\to\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}}$ is the canonical quotient map. Then there exists a commutative diagram
[…]
of equivariant *-epimorphisms such that $\Psi$ is a *-isomorphism.
Proof..
Since ${\mathcal{I}}$ is an induced ideal we have that $q_{\mathcal{I}}$ is equivariant. Moreover $\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}}$ admits a coaction and let $q_{\mathcal{I}}({\mathcal{B}}):=\{[\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}}]_{g}\}_{g\in G}$ be the induced Fell bundle. In order to make a distinction, we will write
$\hat{j}^{{\mathcal{B}}}\colon{\mathcal{B}}\to\mathrm{C}^{*}_{\max}({\mathcal{B}})\quad\text{and}\quad\hat{j}^{q_{{\mathcal{I}}}({\mathcal{B}})}\colon q_{{\mathcal{I}}}({\mathcal{B}})\to\mathrm{C}^{*}_{\max}(q_{\mathcal{I}}({\mathcal{B}}))$
for the induced embeddings of the corresponding Fell bundles.
First note that the family $\left\{\hat{j}^{q_{\mathcal{I}}({\mathcal{B}})}_{g}\circ q_{{\mathcal{I}}}\circ\hat{j}_{g}^{{\mathcal{B}}}\right\}_{g\in G}$ is a representation of ${\mathcal{B}}$ and hence there is an induced *-epimorphism
$\Phi\colon\mathrm{C}^{*}_{\max}({\mathcal{B}})\to\mathrm{C}^{*}_{\max}(q_{\mathcal{I}}({\mathcal{B}})).$
By the definition of $\Phi$ we have that
${\mathcal{I}}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e}={\mathcal{I}}\cap\hat{j}^{{\mathcal{B}}}_{e}({\mathcal{B}}_{e})\subseteq\hat{j}^{{\mathcal{B}}}_{e}\left(\ker\left(\hat{j}^{q_{\mathcal{I}}({\mathcal{B}})}_{e}\circ q_{{\mathcal{I}}}\circ\hat{j}^{{\mathcal{B}}}_{e}\right)\right)\subseteq\ker\Phi,$
and thus we have that
${\mathcal{I}}=\left\langle{\mathcal{I}}\cap[\mathrm{C}^{*}_{\max}({\mathcal{B}})]_{e}\right\rangle\subseteq\ker\Phi.$
Therefore we obtain a commutative diagram
[…]
of *-epimorphisms. By checking on the generators we have that the *-epimorphisms are equivariant.
On the other hand, since $q_{\mathcal{I}}({\mathcal{B}})$ is a topological C*-grading of $\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}}$, by [17, Theorem 19.5] there exists a canonical *-epimorphism
$\mathrm{C}^{*}_{\max}(q_{\mathcal{I}}({\mathcal{B}}))\to\mathrm{C}^{*}_{\max}({\mathcal{B}})/{\mathcal{I}},$
which is the inverse of $\Psi$. Hence $\Psi$ is a *-isomorphism, and the proof is complete. ∎
2.2. Semigroups
We will require some elements on semigroup algebras and right ideals of a semigroup from the work of Li [36], while we also fix notation. For a unital discrete left-cancellative semigroup $P$, we let the left creation (isometric) operators given by
$V_{p}\colon\ell^{2}(P)\longrightarrow\ell^{2}(P);\delta_{s}\mapsto\delta_{ps},$
and we write $\mathrm{C}^{*}_{\lambda}(P):=\mathrm{C}^{*}(V_{p}\mid p\in P)$. The restriction to the diagonal $\ell^{\infty}(P)$ induces a faithful conditional expectation on $\mathrm{C}^{*}_{\lambda}(P)$. For a set $Z\subseteq P$ we will write $E_{[Z]}$ for the projection on [$[\{\delta_{p}\mid p\in Z\}]\subseteq\ell^{2}(P)$. It follows that
$E_{[Z_{1}]}E_{[Z_{2}]}=E_{[Z_{1}\cap Z_{2}]}\text{ for all }Z_{1},Z_{2}\subseteq P.$
For a set $Z\subseteq P$ and $p\in P$ we write
$pZ:=\{px\mid x\in Z\}\quad\text{and}\quad p^{-1}Z:=\{y\in P\mid py\in Z\}.$
By definition we have that $p^{-1}P=P$. We write ${\mathcal{J}}$ for the smallest family of right ideals of $P$ containing $P$ and $\emptyset$ that is closed under left multiplication, taking pre-images under left multiplication (as in the sense above) and finite intersections, i.e.,
${\mathcal{J}}:=\left\{\bigcap\limits_{j=1}^{N}q_{j,n_{j}}^{-1}p_{j,n_{j}}\dots q_{j,1}^{-1}p_{j,1}P\mid N,n_{j}\in\mathbb{N};p_{j,k},q_{j,k}\in P\right\}\bigcup\{\emptyset\}.$
The right ideals of $P$ in ${\mathcal{J}}$ are called constructible. By the proof of [36, Lemma 3.3] we obtain a reduced form for the elements of ${\mathcal{J}}$, since
$q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}p_{1}^{-1}q_{1}\dots p_{n}^{-1}q_{n}Z=(q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P)\cap Z$
for every $p_{i},q_{i}\in P$ and every subset $Z$ of $P$. Consequently the set of constructible ideals is automatically closed under finite intersections, i.e.,
${\mathcal{J}}=\left\{q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P\mid n\in\mathbb{N};p_{i},q_{i}\in P\right\}\cup\{\emptyset\}.$
We will write ${\mathbf{x}},{\mathbf{y}}$ etc. for the elements in ${\mathcal{J}}$.
Henceforth we will assume that $P$ is a unital subsemigroup of a discrete group $G$. In this case we have that
$q^{-1}pZ=P\cap q^{-1}\cdot p\cdot Z,\text{ for }Z\subseteq P,$
and therefore inductively we obtain that
$q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P=P\cap(q_{n}^{-1}\cdot p_{n}\cdot P)\cap(q_{n}^{-1}\cdot p_{n}\cdot q_{n-1}^{-1}\cdot p_{n-1}\cdot P)\cap\cdots\cap(q_{n}^{-1}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}\cdot P).$
Note also that for every ${\mathbf{x}}\in{\mathcal{J}}$ we can pick $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ such that $p_{1}^{-1}q_{1}\cdots p_{n}^{-1}q_{n}=e$ and ${\mathbf{x}}=q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P$. Moreover, $\mathrm{C}^{*}_{\lambda}(P)$ admits a normal coaction by $G$ such that
$[\mathrm{C}^{*}_{\lambda}(P)]_{g}=\overline{\operatorname{span}}\{V_{p_{1}}^{*}V_{q_{1}}\cdots V_{p_{n}}^{*}V_{q_{n}}\mid n\in\mathbb{N};p_{1},q_{1},\dots,p_{n},q_{n}\in P;p_{1}^{-1}q_{1}\cdots p_{n}^{-1}q_{n}=g\}.$
The coaction by $G$ is implemented by the unitary operator
$U\colon\ell^{2}(P)\otimes\ell^{2}(G)\to\ell^{2}(P)\otimes\ell^{2}(G);U(\delta_{s}\otimes\delta_{g})=\delta_{s}\otimes\delta_{sg}\text{ for all }s\in P,g\in G.$
A routine calculation shows that the *-homomorphism
$\mathrm{C}^{*}_{\lambda}(P)\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\mathrm{C}^{*}(V_{p}\otimes I\mid p\in P)\stackrel{{\scriptstyle\operatorname{ad}_{U}}}{{\longrightarrow}}\mathrm{C}^{*}(V_{p}\otimes\lambda_{p}\mid p\in P)$
is a reduced coaction, and thus it lifts to a normal coaction on $\mathrm{C}^{*}_{\lambda}(P)$. The induced faithful conditional expectation on $\mathrm{C}^{*}_{\lambda}(P)$ coincides with the compression to the diagonal $\ell^{\infty}(P)$. In the presence of an overlying group there is an explicit formula for the projection $E_{[{\mathbf{x}}]}$ for ${\mathbf{x}}\in{\mathcal{J}}$, obtained in [36, Lemma 3.1], i.e.,
$E_{[{\mathbf{x}}]}=V_{p_{1}}^{*}V_{q_{1}}\cdots V_{p_{n}}^{*}V_{q_{n}},$
for any $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ satisfying ${\mathbf{x}}=q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P$ and $p_{1}^{-1}q_{1}\cdots p_{n}^{-1}q_{n}=e$ in $G$.
2.3. C*-correspondences
The theory of Hilbert modules over C*-algebras is well-developed. The reader is addressed to [35, 39] for an excellent introduction to the subject.
A C*-correspondence $X$ over $A$ is a right Hilbert module over $A$ with a left action given by a *-homomorphism $\varphi_{X}\colon A\to{\mathcal{L}}(X)$, where ${\mathcal{L}}(X)$ denotes the C*-algebra of adjointable operators on $X$. A C*-correspondence $X$ over $A$ is called non-degenerate if $[\varphi_{X}(A)X]=X$. We write ${\mathcal{K}}(X)$ for the closed linear span of the rank one adjointable operators $\theta_{\xi,\eta}$. For two C*-correspondences $X,Y$ over the same $A$ we write $X\otimes_{A}Y$ for the balanced tensor product over $A$. We say that $X$ is unitarily equivalent to $Y$ (symb. $X\simeq Y$) if there is a surjective adjointable operator $U\in{\mathcal{L}}(X,Y)$ that is an $A$-bimodule map and $\left\langle U\xi,U\eta\right\rangle=\left\langle\xi,\eta\right\rangle$ for all $\xi,\eta\in X$.
A representation of a C*-correspondence $X$ over $A$ is a pair $(t_{0},t_{1})$ where $t_{0}\colon A\to{\mathcal{B}}(H)$ is a *-homomorphism and $t_{1}\colon X\to{\mathcal{B}}(H)$ is a linear map that satisfies the following:
  1. (i)
    $t_{0}(a)t_{1}(\xi)=t_{1}(\varphi_{X}(a)\xi)$ for all $\xi\in X$ and $a\in A$,
  2. (ii)
    $t_{1}(\xi)^{*}t_{1}(\eta)=t_{0}(\left\langle\xi,\eta\right\rangle)$ for all $\xi,\eta\in X$.
We note that condition (ii) also implies that $t_{1}(\xi)t_{0}(a)=t_{1}(\xi a)$. Every representation $(t_{0},t_{1})$ as above defines a *-representation
$t^{(1)}\colon{\mathcal{K}}(X)\to{\mathcal{B}}(H);\theta_{\xi,\eta}\mapsto t(\xi)t(\eta)^{*}\text{ for all }\xi,\eta\in X.$
If $t_{0}$ is injective, then both $t_{1}$ and $t^{(1)}$ are isometric.
2.4. Product systems
Let $P$ be a unital left-cancellative discrete semigroup. We say that a family $X=\{X_{p}\}_{p\in P}$ of closed operator spaces in a common ${\mathcal{B}}(H)$ is a (concrete) product system if the following are satisfied:
  1. (i)
    $A:=X_{e}$ is a C*-algebra;
  2. (ii)
    $X_{p}\cdot X_{q}\subseteq X_{pq}$ for all $p,q\in P$;
  3. (iii)
    $X_{p}^{*}\cdot X_{pq}\subseteq X_{q}$ for all $p,q\in P$.
Uniqueness of $q\in P$ in item (iii) follows since $P$ is left-cancellative. Moreover it follows that each $X_{p}$ is a C*-correspondence over $A$.
The properties of a concrete product system are enough to provide a Fock space representation. Towards this end, consider $X$ in some ${\mathcal{B}}(H)$, and let ${\mathcal{B}}(H)$ with its trivial C*-correspondence structure. We will be writing
$\left\langle\cdot,\cdot\right\rangle_{p}\colon X_{p}\times X_{p}\longrightarrow A;(\xi_{p},\eta_{p})\mapsto\left\langle\xi_{p},\eta_{p}\right\rangle_{p}:=\xi_{p}^{*}\cdot\eta_{p},$
for the inner product induced on the $X_{p}$. For every $\xi_{p}\in X_{p}$ we define the multiplication operator
$M_{\xi_{p}}^{q,pq}\colon X_{q}\longrightarrow X_{pq};\eta_{q}\mapsto\xi_{p}\cdot\eta_{q},$
with the multiplication taking place inside ${\mathcal{B}}(H)$, satisfying
$\|M_{\xi_{p}}^{q,pq}\|_{{\mathcal{B}}(X_{q},X_{pq})}\leq\left\|\xi_{p}\right\|_{X_{p}}.$
Associativity of the product gives that $M_{\xi_{p}}^{q,pq}\in{\mathcal{L}}(X_{q},X_{pq})$ with
$(M_{\xi_{p}}^{q,pq})^{*}\colon X_{pq}\to X_{q};\eta_{pq}\mapsto\xi_{p}^{*}\cdot\eta_{pq}\in X_{q}.$
Consider the Fock space ${\mathcal{F}}X:=\sum^{\oplus}_{r\in P}X_{r}$, as a right Hilbert module over $A$. For every $\xi_{p}\in X_{p}$, define the left creation operator
$\lambda_{p}(\xi_{p}):=\operatornamewithlimits{\sum^{\oplus}}_{r\in P}M_{\xi_{p}}^{r,pr}\;\text{ so that }\;\lambda_{p}(\xi_{p})^{*}=\operatornamewithlimits{\sum^{\oplus}}_{r\in P}(M_{\xi_{p}}^{r,pr})^{*},$
where the sum is taken in the s*-topology, and with the understanding that we are embedding ${\mathcal{L}}(X_{q},X_{pq})\hookrightarrow{\mathcal{L}}({\mathcal{F}}X)$ as the $(pq,q)$-entry. By applying on $\eta_{r}\in X_{r}\subseteq{\mathcal{F}}X$ we derive that
$\lambda_{p}(\xi_{p})\eta_{r}=\xi_{p}\cdot\eta_{r}\quad\text{and}\quad\lambda_{p}(\xi_{p})^{*}\eta_{r}=\begin{cases}\xi_{p}^{*}\cdot\eta_{r}&\text{if }r\in pP,\\ 0&\text{if }r\notin pP.\end{cases}$
We write $\lambda\colon X\to{\mathcal{L}}({\mathcal{F}}X)$ for this map, and we refer to $\lambda=\{\lambda_{p}\}_{p\in P}$ as the Fock representation of $X$. We will write ${\mathcal{T}}_{\lambda}(X)$ for the Fock C*-algebra defined as $\mathrm{C}^{*}(\lambda_{p}(X_{p})\mid p\in P)$.
More generally, a (Toeplitz) representation $t=\{t_{p}\}_{p\in P}$ of $X$ consists of a family of linear maps $t_{p}$ of $X_{p}$ such that:
  1. (i)
    $t_{e}$ is a *-representation of $A:=X_{e}$;
  2. (ii)
    $t_{p}(\xi_{p})t_{q}(\xi_{q})=t_{pq}(\xi_{p}\xi_{q})$ for all $\xi_{p}\in X_{p}$ and $\xi_{q}\in X_{q}$;
  3. (iii)
    $t_{p}(\xi_{p})^{*}t_{pq}(\xi_{pq})=t_{q}(\xi_{p}^{*}\xi_{pq})$ for all $\xi_{p}\in X_{p}$ and $\xi_{pq}\in X_{pq}$.
By definition every pair $(t_{e},t_{p})$ is a representation of the C*-correspondence $X_{p}$ over $A$. A representation $t$ is called injective if $t_{e}$ is injective on $A$. It transpires that the Fock representation is an injective representation of $X$. The Toeplitz algebra ${\mathcal{T}}(X)$ of $X$ is the universal C*-algebra generated by $X=\{X_{p}\}_{p\in P}$ with respect to the representations of $X$. If $t$ defines a representation of $X$, then we will write $t_{\ast}\colon{\mathcal{T}}(X)\to{\mathcal{B}}(H)$ for the induced *-representation of ${\mathcal{T}}(X)$ and
$\mathrm{C}^{*}(t)\equiv\mathrm{C}^{*}(t_{\ast}):=t_{\ast}({\mathcal{T}}(X))=\mathrm{C}^{*}(t_{p}(X_{p})\mid p\in P).$
Henceforth we will assume that $P$ is a unital subsemigroup of a group $G$.
Proposition 2.3.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $t$ be a representation of $X$ and let $\xi_{r}\in X_{r}$, $\xi_{p_{i}}\in X_{p_{i}}$, and $\xi_{q_{i}}\in X_{q_{i}}$ for $r,p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. If $r\in q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$, then $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r\in P$ and
$\left(t_{p_{1}}(\xi_{p_{1}})^{*}\right)^{\varepsilon}t_{q_{1}}(\xi_{q_{1}})\cdots t_{p_{n}}(\xi_{p_{n}})^{*}t_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}t_{r}(\xi_{r})=t_{p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r}\left((\xi_{p_{1}}^{*})^{\varepsilon}\xi_{q_{1}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}\right).$
Proof..
We start with the following three comments. First, suppose that $s=p^{-1}r$ for some $p,s\in P$. Then by definition we have that
$t_{p}(\xi_{p})^{*}t_{r}(\xi_{r})=t_{p}(\xi_{p})^{*}t_{ps}(\xi_{r})=t_{s}(\xi_{p}^{*}\xi_{r}).$
Next, suppose that $s=qr$ for some $q,s\in P$. Then by definition we have that
$t_{q}(\xi_{q})t_{r}(\xi_{r})=t_{qr}(\xi_{q}\xi_{r})=t_{s}(\xi_{q}\xi_{r}).$
Finally, suppose that $s=p^{-1}qr$ for some $p,q,s\in P$. Then we have that
$t_{p}(\xi_{p})^{*}t_{q}(\xi_{q})t_{r}(\xi_{r})=t_{p}(\xi_{p})^{*}t_{qr}(\xi_{q}\xi_{r})=t_{p}(\xi_{p})^{*}t_{ps}(\xi_{q}\xi_{r})=t_{s}(\xi_{p}^{*}\xi_{q}\xi_{r}).$
We proceed with the proof. Let $r$ be in $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ for some $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. We have that there is an $s_{1}\in q_{n-1}^{-1}p_{n-1}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ such that $s_{1}=p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r$. From the comments above we then have that $\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}\in X_{s_{1}}$ and
$t_{p_{n}}(\xi_{p_{n}})^{*}t_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}t_{r}(\xi_{r})=t_{s_{1}}(\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}).$
Since $s_{1}\in q_{n-1}^{-1}p_{n-1}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ we have that there is an $s_{2}\in q_{n-2}^{-1}p_{n-2}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ such that $s_{2}=p_{n-1}^{-1}q_{n-1}s_{1}$, and therefore
$t_{p_{n-1}}(\xi_{p_{n-1}})^{*}t_{q_{n-1}}(\xi_{q_{n-1}})t_{s_{1}}(\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r})=t_{s_{2}}(\xi_{p_{n-1}}^{*}\xi_{q_{n-1}}\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}).$
Continuing inductively for each $k=3,\dots,n-1$ we obtain an $s_{k}\in q_{n-k}^{-1}p_{n-k}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ such that $s_{k}=p_{n-(k-1)}^{-1}q_{n-(k-1)}s_{k-1}$, and for each $k=1,\dots,n-1$ we have that
$t_{p_{n-(k-1)}}(\xi_{p_{n-(k-1)}})^{*}t_{q_{n-(k-1)}}(\xi_{q_{n-(k-1)}})\cdots t_{p_{n}}(\xi_{p_{n}})^{*}t_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}t_{r}(\xi_{r})=$
$\hskip 142.26378pt=t_{s_{k}}(\xi_{p_{n-(k-1)}}^{*}\xi_{q_{n-(k-1)}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r})$
$\hskip 142.26378pt=t_{p_{n-(k-1)}^{-1}q_{n-(k-1)}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r}(\xi_{p_{n-(k-1)}}^{*}\xi_{q_{n-(k-1)}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}).$
Note that $r=q_{n}^{-\varepsilon^{\prime}}p_{n}\cdots q_{1}^{-1}p_{1}^{\varepsilon}s_{n}$ for some $s_{n}\in P$, and thus $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r=s_{n}\in P$. Since $p_{2}^{-1}q_{2}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r=q_{1}^{-1}p_{1}^{\varepsilon}s_{n}$, from the comments above we obtain
$(t_{p_{1}}(\xi_{p_{1}})^{*})^{\varepsilon}t_{q_{1}}(\xi_{q_{1}})t_{p_{2}}(\xi_{p_{2}})^{*}t_{q_{2}}(\xi_{q_{2}})\cdots t_{p_{n}}(\xi_{p_{n}})^{*}t_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}t_{r}(\xi_{r})=$
$\hskip 142.26378pt=(t_{p_{1}}(\xi_{p_{1}})^{*})^{\varepsilon}t_{q_{1}}(\xi_{q_{1}})t_{p_{2}^{-1}q_{2}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r}(\xi_{p_{2}}^{*}\xi_{q_{2}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r})$
$\hskip 142.26378pt=(t_{p_{1}}(\xi_{p_{1}})^{*})^{\varepsilon}t_{q_{1}}(\xi_{q_{1}})t_{q_{1}^{-1}p_{1}^{\varepsilon}s_{n}}(\xi_{p_{2}}^{*}\xi_{q_{2}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r})$
$\hskip 142.26378pt=t_{s_{n}}((\xi_{p_{1}}^{*})^{\varepsilon}\xi_{q_{1}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}),$
as required. ∎
For the Fock representation we also have the following proposition.
Proposition 2.4.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$, and $\xi_{p_{i}}\in X_{p_{i}}$ and $\xi_{q_{i}}\in X_{q_{i}}$ for $i=1,\dots,n$. Then for each $r\in P$ and $\xi_{r}\in X_{r}$ we have
$\left(\lambda_{p_{1}}(\xi_{p_{1}})^{*}\right)^{\varepsilon}\lambda_{q_{1}}(\xi_{q_{1}})\cdots\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\xi_{r}=\begin{cases}(\xi_{p_{1}}^{*})^{\varepsilon}\xi_{q_{1}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}&\text{if }r\in q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P,\\ 0&\text{if }r\not\in q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P.\end{cases}$
Proof..
If $r\in q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $a\in A:=X_{e}$, then by Proposition 2.3 we have that
$\left(\lambda_{p_{1}}(\xi_{p_{1}})^{*}\right)^{\varepsilon}\lambda_{q_{1}}(\xi_{q_{1}})\cdots\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\xi_{r}a=$
$\hskip 113.81102pt=\left(\lambda_{p_{1}}(\xi_{p_{1}})^{*}\right)^{\varepsilon}\lambda_{q_{1}}(\xi_{q_{1}})\cdots\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\lambda_{r}(\xi_{r})a$
$\hskip 113.81102pt=\lambda_{p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}r}((\xi_{p_{1}}^{*})^{\varepsilon}\xi_{q_{1}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r})a$
$\hskip 113.81102pt=(\xi_{p_{1}}^{*})^{\varepsilon}\xi_{q_{1}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}a.$
Since $[X_{r}A]=X_{r}$ we get
$(\lambda_{p_{1}}(\xi_{p_{1}})^{*})^{\varepsilon}\lambda_{q_{1}}(\xi_{q_{1}})\cdots\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\xi_{r}=(\xi_{p_{1}}^{*})^{\varepsilon}\xi_{q_{1}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}.$
Next let $r\not\in q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$, and suppose towards a contradiction that
$(\lambda_{p_{1}}(\xi_{p_{1}})^{*})^{\varepsilon}\lambda_{q_{1}}(\xi_{q_{1}})\cdots\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\xi_{r}\neq 0.$
In particular we have that
$\lambda_{p_{n}}(\xi_{p_{n}})^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}=\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\xi_{r}\neq 0.$
Hence $p_{n}r_{1}=q_{n}^{\varepsilon^{\prime}}r$ for some $r_{1}\in P$ and
$0\neq\lambda_{p_{n}}(\xi_{p_{n}})^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}=\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}\in X_{r_{1}}.$
Inductively, for $k=1,\dots,n$ we obtain $r_{k}\in P$ such that
$p_{n}r_{1}=q_{n}^{\varepsilon^{\prime}}r,\,p_{1}^{\varepsilon}r_{n}=q_{1}r_{n-1},\text{ and }p_{n-(k-1)}r_{k}=q_{n-(k-1)}r_{k-1}\text{ for all}k=2,\dots,n-1,$
and therefore $r=q_{n}^{-\varepsilon^{\prime}}p_{n}\cdots q_{n-(k-1)}^{-1}p_{n-(k-1)}r_{k}$ for each $k=2,\dots,n$. Thus,
$r\in P\cap(q_{n}^{-\varepsilon^{\prime}}\cdot p_{n}\cdot P)\cap(q_{n}^{-\varepsilon^{\prime}}\cdot p_{n}\cdot q_{n-1}^{-1}\cdot p_{n-1}\cdot P)\cap\cdots\cap(q_{n}^{-\varepsilon^{\prime}}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}^{\varepsilon}\cdot P)=q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P,$
which is a contradiction. ∎
A representation $t=\{t_{p}\}_{p\in P}$ of $X$ will be called equivariant if there exists a *-homomorphism $\delta$ of $\mathrm{C}^{*}(t)$ such that
$\delta\colon\mathrm{C}^{*}(t)\to\mathrm{C}^{*}(t)\otimes\mathrm{C}^{*}_{\max}(G);t_{p}(\xi_{p})\mapsto t_{p}(\xi_{p})\otimes u_{p}.$
It follows that $\delta$ is injective with a left inverse given by the map ${\operatorname{id}}\otimes\chi$. Moreover, it satisfies the coaction identity and hence $\mathrm{C}^{*}(t)$ admits a coaction by $G$. For simplicity, we will say that $t$ admits a coaction by $G$ if such a $\delta$ exists. In this case the $g$-fiber $[\mathrm{C}^{*}(t)]_{g}$, for $g\in G$, is the closed linear span of the elements
$\left(t_{p_{1}}(X_{p_{1}})^{*}\right)^{\varepsilon}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})^{\varepsilon^{\prime}}\text{ such that }p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=g,$
for $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ and $n\in\mathbb{N}$. A proof can be found in [47, Lemma 2.2] for non-degenerate product systems, but similar arguments give the conclusion in the general case.
Let $t$ be a representation of $X$. For ${\mathbf{x}}\in{\mathcal{J}}$ we define the $\mathbf{K}$-core on ${\mathbf{x}}$ of $\mathrm{C}^{*}(t)$ to be the closed linear span of the spaces
$\left(t_{p_{1}}(X_{p_{1}})^{*}\right)^{\varepsilon}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})^{\varepsilon^{\prime}}$
for any $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ that satisfy
$p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e\text{ and }q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P={\mathbf{x}}.$
We write $\mathbf{K}_{{\mathbf{x}},t_{\ast}}$ for this closed space. We do not claim that $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$. Note that, if $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$, then
$q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P=p_{1}^{-\varepsilon}q_{1}\dots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}P.$
Indeed, since $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ we have that
$q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ $=P\cap(q_{n}^{-\varepsilon^{\prime}}\cdot p_{n}\cdot P)\cap\cdots\cap(q_{n}^{-\varepsilon^{\prime}}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}^{\varepsilon}\cdot P)$
$=(p_{1}^{-\varepsilon}\cdot q_{1}\cdots p_{n}^{-1}\cdot q_{n}^{\varepsilon^{\prime}}\cdot P)\cap(p_{1}^{-\varepsilon}\cdot q_{1}\cdots p_{n-1}^{-1}\cdot q_{n-1}\cdot P)\cap\cdots\cap P$
$=p_{1}^{-\varepsilon}q_{1}\dots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}P.$
We conclude that each $\mathbf{K}_{{\mathbf{x}},t_{\ast}}$ is a selfadjoint space. Moreover, if we have
${\mathbf{x}}=q_{n}^{-\varepsilon_{1}^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon_{1}}P\quad\text{and}\quad{\mathbf{y}}=s_{m}^{-\varepsilon_{2}^{\prime}}r_{m}\dots s_{1}^{-1}r_{1}^{\varepsilon_{2}}P,$
with $p_{1}^{-\varepsilon_{1}}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon_{1}^{\prime}}=e$ and $r_{1}^{-\varepsilon_{2}}s_{1}\cdots r_{m}^{-1}s_{m}^{\varepsilon_{2}^{\prime}}=e$, then
${\mathbf{x}}\cap{\mathbf{y}}$ $=\left(P\cap(q_{n}^{-\varepsilon_{1}^{\prime}}\cdot p_{n}\cdot P)\cap\cdots\cap(q_{n}^{-\varepsilon_{1}^{\prime}}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}^{\varepsilon_{1}}\cdot P)\right)\bigcap$
$\hskip 85.35826pt\bigcap\left(P\cap(s_{m}^{-\varepsilon_{2}^{\prime}}\cdot r_{m}\cdot P)\cap\cdots\cap(s_{m}^{-\varepsilon_{2}^{\prime}}\cdot r_{m}\cdots s_{1}^{-1}\cdot r_{1}^{\varepsilon_{2}}\cdot P)\right)$
$=P\cap(q_{n}^{-\varepsilon_{1}^{\prime}}\cdot p_{n}\cdot P)\cap\cdots\cap(q_{n}^{-\varepsilon_{1}^{\prime}}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}^{\varepsilon_{1}}\cdot P)\bigcap$
$\hskip 85.35826pt\bigcap(q_{n}^{-\varepsilon_{1}^{\prime}}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}^{\varepsilon_{1}}\cdot s_{m}^{-\varepsilon_{2}^{\prime}}\cdot r_{m}\cdot P)\cap\cdots$
$\hskip 113.81102pt\cdots\cap(q_{n}^{-\varepsilon_{1}^{\prime}}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}^{\varepsilon_{1}}\cdot s_{m}^{-\varepsilon_{2}^{\prime}}\cdot r_{m}\cdots s_{1}^{-1}\cdot r_{1}^{\varepsilon_{2}}\cdot P)$
$=q_{n}^{-\varepsilon_{1}^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon_{1}}s_{m}^{-\varepsilon_{2}^{\prime}}r_{m}\dots s_{1}^{-1}r_{1}^{\varepsilon_{2}}P,$
with $r_{1}^{-\varepsilon_{2}}s_{1}\cdots r_{m}^{-1}s_{m}^{\varepsilon_{2}^{\prime}}p_{1}^{-\varepsilon_{1}}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon_{1}^{\prime}}=e$. Hence we obtain that
$\mathbf{K}_{{\mathbf{x}},t_{\ast}}\cdot\mathbf{K}_{{\mathbf{y}},t_{\ast}}\subseteq\mathbf{K}_{{\mathbf{x}}\cap{\mathbf{y}},t_{\ast}}.$
From this we derive that every $\mathbf{K}_{{\mathbf{x}},t_{\ast}}$ is an algebra (as ${\mathbf{x}}\cap{\mathbf{x}}={\mathbf{x}}$), and thus a C*-subalgebra of $\mathrm{C}^{*}(t)$. More generally, for a finite $\cap$-closed ${\mathcal{F}}\subseteq{\mathcal{J}}$ we define the $\mathbf{B}$-core on ${\mathcal{F}}$ by
$\mathbf{B}_{{\mathcal{F}},t_{\ast}}:=\sum_{{\mathbf{x}}\in{\mathcal{F}}}\mathbf{K}_{{\mathbf{x}},t_{\ast}}.$
By the discussion above, it follows that if ${\mathcal{F}}$ is $\cap$-closed and ${\mathbf{x}}\in{\mathcal{J}}$ is such that ${\mathbf{x}}\cap{\mathbf{y}}\in{\mathcal{F}}$ for every ${\mathbf{y}}\in{\mathcal{F}}$, then $\mathbf{B}_{{\mathcal{F}},t_{\ast}}$ is an ideal in $\mathbf{B}_{{\mathcal{F}}\cup\{{\mathbf{x}}\},t_{\ast}}$. An induction argument gives that $\mathbf{B}_{{\mathcal{F}},t_{\ast}}$ is a C*-subalgebra of $\mathrm{C}^{*}(t)$ when ${\mathcal{F}}\subseteq{\mathcal{J}}$ is finite and $\cap$-closed. In particular we have that, if $t$ is equivariant, then the fixed point algebra $[\mathrm{C}^{*}(t)]_{e}$ is the inductive limit of the C*-subalgebras $\mathbf{B}_{{\mathcal{F}},t_{\ast}}$ over $\cap$-closed and finite ${\mathcal{F}}\subseteq{\mathcal{J}}$ (with respect to inclusion of sets).
Proposition 2.5.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $t$ be a representation of $X$ and let $\hat{t}$ be the representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. For $\emptyset\neq{\mathbf{x}}\in{\mathcal{J}}$ and $r\in{\mathbf{x}}$ we have that
$t_{\ast}(b_{{\mathbf{x}}})t_{r}(\xi_{r})=t_{r}(\lambda_{\ast}(b_{{\mathbf{x}}})\xi_{r})\text{ for all }b_{{\mathbf{x}}}\in\mathbf{K}_{{\mathbf{x}},\hat{t}_{\ast}},\xi_{r}\in X_{r}.$
Proof..
First consider an element of the form
$b_{\mathbf{x}}:=\left({\hat{t}_{p_{1}}(\xi_{p_{1}})}^{*}\right)^{\varepsilon}{\hat{t}_{q_{1}}(\xi_{q_{1}})}\cdots{\hat{t}_{p_{n}}(\xi_{p_{n}})}^{*}\hat{t}_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\in\mathbf{K}_{\mathbf{x},\hat{t}_{\ast}},$
such that ${\mathbf{x}}=q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ for some $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. Let $r\in{\mathbf{x}}$. By Proposition 2.3 and Proposition 2.4 we have that
$t_{\ast}(b_{{\mathbf{x}}})t_{r}(\xi_{r})$ $=t_{\ast}\left(\left({\hat{t}_{p_{1}}(\xi_{p_{1}})}^{*}\right)^{\varepsilon}{\hat{t}_{q_{1}}(\xi_{q_{1}})}\cdots{\hat{t}_{p_{n}}(\xi_{p_{n}})}^{*}\hat{t}_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\right)t_{r}(\xi_{r})$
$=\left(t_{p_{1}}(\xi_{p_{1}})^{*}\right)^{\varepsilon}t_{q_{1}}(\xi_{q_{1}})\cdots t_{p_{n}}(\xi_{p_{n}})^{*}t_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}t_{r}(\xi_{r})$
$=t_{r}\left((\xi_{p_{1}}^{*})^{\varepsilon}\xi_{q_{1}}\cdots\xi_{p_{n}}^{*}\xi_{q_{n}}^{\varepsilon^{\prime}}\xi_{r}\right)=t_{r}(\lambda_{\ast}(b_{\mathbf{x}})\xi_{r}).$
Taking finite linear combinations and their norm-limits completes the proof. ∎
The following proposition is an immediate consequence of Proposition 2.4.
Proposition 2.6.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\hat{t}$ be the representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. Then for every ${\mathbf{x}}\in{\mathcal{J}}$ and $r\not\in{\mathbf{x}}$ we have that
$\lambda_{\ast}(b_{\mathbf{x}})\xi_{r}=0\text{ for all }b_{\mathbf{x}}\in\mathbf{K}_{\mathbf{x},\hat{t}_{\ast}},\xi_{r}\in X_{r}.$
Proof..
We consider an element of the form
$b_{\mathbf{x}}:=\left({\hat{t}_{p_{1}}(\xi_{p_{1}})}^{*}\right)^{\varepsilon}{\hat{t}_{q_{1}}(\xi_{q_{1}})}\cdots{\hat{t}_{p_{n}}(\xi_{p_{n}})}^{*}\hat{t}_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\in\mathbf{K}_{\mathbf{x},\hat{t}_{\ast}},$
such that ${\mathbf{x}}=q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P$ and $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ for some $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$. Let $\xi_{r}\in X_{r}$. By Proposition 2.4 we obtain that
$\lambda_{\ast}(b_{\mathbf{x}})\xi_{r}=\left(\lambda_{p_{1}}(\xi_{p_{1}})^{*}\right)^{\varepsilon}\lambda_{q_{1}}(\xi_{q_{1}})\cdots\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}\xi_{r}=0.$
Taking finite linear combinations and their norm-limits completes the proof. ∎
By the universal property, the Toeplitz C*-algebra ${\mathcal{T}}(X)$ admits a coaction by $G$. We let
${\mathcal{P}}{\mathcal{S}}X:=\Big{\{}[{\mathcal{T}}(X)]_{g}\Big{\}}_{g\in G}$
be the induced Fell bundle. Moreover, the embedding $X\hookrightarrow{\mathcal{T}}(X)$ is injective, since the Fock representation $\lambda$ is injective. Therefore for each $p\in P$ we have an embedding $X_{p}\hookrightarrow[{\mathcal{P}}{\mathcal{S}}X]_{p}$.
Proposition 2.7.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Then the family $\{X_{p}\hookrightarrow[{\mathcal{P}}{\mathcal{S}}X]_{p}\}_{p\in P}$ of embeddings lifts to a *-isomorphism
${\mathcal{T}}(X)\simeq\mathrm{C}^{*}_{\max}({\mathcal{P}}{\mathcal{S}}X).$
Proof..
Let $\hat{t}$ be the representation of $X$ such that $\mathrm{C}^{*}(\hat{t})={\mathcal{T}}(X)$ and let $\hat{j}\colon{\mathcal{P}}{\mathcal{S}}X\to\mathrm{C}^{*}_{\max}({\mathcal{P}}{\mathcal{S}}X)$ be the induced canonical embedding of the Fell bundle. Let $t=\{t_{p}\}_{p\in P}$ be the family of maps defined by $t_{p}\colon X_{p}\to\mathrm{C}^{*}_{\max}({\mathcal{P}}{\mathcal{S}}X)$ such that $t_{p}(\xi_{p})=\hat{j}_{p}(\hat{t}_{p}(\xi_{p}))$ for each $\xi_{p}\in X_{p}$ and $p\in P$. Then $t$ is a representation of $X$. Indeed, for the first axiom, since $\hat{t}_{e}$ and $\hat{j}_{e}$ are *-representations we obtain that $t_{e}$ is a *-representation of $A:=X_{e}$. For the second axiom, let $\xi_{p}\in X_{p}$ and $\xi_{q}\in X_{q}$, then
$t_{p}(\xi_{p})t_{q}(\xi_{q})=\hat{j}_{p}(\hat{t}_{p}(\xi_{p}))\hat{j}_{q}(\hat{t}_{q}(\xi_{q}))=\hat{j}_{pq}(\hat{t}_{pq}(\xi_{p}\xi_{q}))=t_{pq}(\xi_{p}\xi_{q}).$
For the third axiom, let $\xi_{p}\in X_{p}$ and $\xi_{pq}\in X_{pq}$, then
$t_{p}(\xi_{p})^{*}t_{pq}(\xi_{pq})=\hat{j}_{p}(\hat{t}_{p}(\xi_{p}))^{*}\hat{j}_{pq}(\hat{t}_{pq}(\xi_{pq}))=\hat{j}_{p^{-1}}(\hat{t}_{p}(\xi_{p})^{*})\hat{j}_{pq}(\hat{t}_{pq}(\xi_{pq}))=\hat{j}_{q}(\hat{t}_{q}(\xi_{p}^{*}\xi_{pq}))=t_{q}(\xi_{p}^{*}\xi_{pq}),$
as required.
Therefore, there is an induced *-epimorphism
${\mathcal{T}}(X)\to\mathrm{C}^{*}(t)=\mathrm{C}^{*}_{\max}({\mathcal{P}}{\mathcal{S}}X);\hat{t}_{p}(\xi_{p})\mapsto t_{p}(\xi_{p})=\hat{j}_{p}(\hat{t}_{p}(\xi_{p})).$
On the other hand, by [17, Theorem 19.5] we obtain a *-epimorphism
$\mathrm{C}^{*}_{\max}({\mathcal{P}}{\mathcal{S}}X)\to{\mathcal{T}}(X);\hat{j}_{p}(\hat{t}_{p}(\xi_{p}))\mapsto\hat{t}_{p}(\xi_{p})$
in the other direction, and the proof is complete. ∎
As noted in [11, Proposition 4.1], the representation $\lambda$ admits a reduced coaction by using the unitary
$U\colon{\mathcal{F}}X\otimes\ell^{2}(G)\to{\mathcal{F}}X\otimes\ell^{2}(G);U(\xi_{r}\otimes\delta_{g})=\xi_{r}\otimes\delta_{rg}.$
Indeed, for this $U$ we have that
$U\cdot(\lambda_{p}(\xi_{p})\otimes I)=(\lambda_{p}(\xi_{p})\otimes\lambda_{p})\cdot U\text{ for all }p\in P,$
and therefore the map
${\mathcal{T}}_{\lambda}(X)\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\mathrm{C}^{*}(\lambda_{p}(\xi_{p})\otimes I\mid p\in P)\stackrel{{\scriptstyle\operatorname{ad}_{U}}}{{\longrightarrow}}\mathrm{C}^{*}(\lambda_{p}(\xi_{p})\otimes\lambda_{p}\mid p\in P)$
defines a reduced coaction on ${\mathcal{T}}_{\lambda}(X)$. Consequently, it lifts to a normal coaction $\delta$ on ${\mathcal{T}}_{\lambda}(X)$.
2.5. Fowler’s product systems
Fowler [18] defines a product system $X$ over $P$ with coefficients in a C*-algebra $A$ as a family $\{X_{p}\}_{p\in P}$ of C*-correspondences over $A$ together with a family of unitary maps $\{u_{p,q}\colon X_{p}\otimes_{A}X_{q}\to X_{pq}\}_{p,q\in P}$ such that:
  1. (i)
    the space $X_{e}$ is the C*-correspondence $A$ over $A$ where the left and right actions of $A$ is multiplication on $A$;
  2. (ii)
    if $p=e$, then $u_{e,q}\colon A\otimes_{A}X_{q}\to X_{q}$ is given by the left action of $A$ on $X_{q}$ for $q\in P$;
  3. (iii)
    if $q=e$, then $u_{p,e}\colon X_{p}\otimes_{A}A\to X_{p}$ is given by the right action of $A$ on $X_{p}$ for $p\in P$;
  4. (iv)
    the maps $\{u_{p,q}\}_{p,q\in P}$ are associative in the sense that
    $u_{p,qr}\circ({\operatorname{id}}_{X_{p}}\otimes u_{q,r})=u_{pq,r}\circ(u_{p,q}\otimes{\operatorname{id}}_{X_{r}})\text{ for all }p,q,r\in P.$
Note here that item (ii) implies in particular that every $X_{p}$ is non-degenerate. To allow for comparisons, we will refer to $(X,\{u_{p,q}\}_{p,q\in P})$ as a product system in the sense of Fowler.
It can be directly verified that, if $X$ is a concrete product system in some ${\mathcal{B}}(H)$ such that $[X_{p}X_{q}]=X_{pq}$ for all $p,q\in P$, then the family $\{X_{p}\}_{p\in P}$ defines a product system in the sense of Fowler by considering the unitary maps $\{u_{p,q}\}_{p,q\in P}$ given by
$u_{p,q}\colon X_{p}\otimes_{A}X_{q}\to X_{pq};\xi_{p}\otimes\xi_{q}\mapsto\xi_{p}\cdot\xi_{q}.$
A (Toeplitz) representation $t$ of a product system $X$ in the sense of Fowler consists of a family $\{t_{p}\}_{p\in P}$, where $(t_{e},t_{p})$ is a representation of the C*-correspondence $X_{p}$ for all $p\in P$, and
$t_{p}(\xi_{p})t_{q}(\xi_{q})=t_{pq}(u_{p,q}(\xi_{p}\otimes\xi_{q}))\text{ for all }\xi_{p}\in X_{p},\xi_{q}\in X_{q},p,q\in P.$
If $(X,\{u_{p,q}^{X}\}_{p,q\in P})$ and $(Y,\{u_{p,q}^{Y}\}_{p,q\in P})$ are two product systems in the sense of Fowler with coefficients in two C*-algebras $A$ and $B$ respectively, then we say that $X$ is unitarily equivalent to $Y$ if there is a family $\{W_{p}\colon X_{p}\to Y_{p}\}_{p\in P}$ of unitaries such that:
  1. (i)
    $W_{e}\colon A\to B$ is a *-isomorphism;
  2. (ii)
    $\left\langle W_{p}(\xi_{p}),W_{p}(\eta_{p})\right\rangle_{Y_{p}}=W_{e}(\left\langle\xi_{p},\eta_{p}\right\rangle_{X_{p}})$ for all $\xi_{p},\eta_{p}\in X_{p}$ and $p\in P\setminus\{e\}$;
  3. (iii)
    $\varphi_{Y_{p}}(W_{e}(a))W_{p}(\xi_{p})=W_{p}(\varphi_{X_{p}}(a)\xi_{p})$ for all $a\in A,\xi_{p}\in X_{p}$ and $p\in P\setminus\{e\}$;
  4. (iv)
    $W_{p}(\xi_{p})W_{e}(a)=W_{p}(\xi_{p}a)$ for all $a\in A,\xi_{p}\in X_{p}$ and $p\in P\setminus\{e\}$;
  5. (v)
    $u_{p,q}^{Y}\circ(W_{p}\otimes W_{q})=W_{pq}\circ u_{p,q}^{X}$ for all $p,q\in P$.
In this case the representations of $X$ are in bijection with the representations of $Y$. The reader may refer to [9] for the full details.
Given a product system $X$ in the sense of Fowler, we can define the left-creation operators
$\lambda_{p}(\xi_{p}):=\sum_{r\in P}\tau_{r}^{pr}(\xi_{p})\text{ for all }\xi_{p}\in X_{p},$
as the s*-sum of the adjointable operators
$\tau_{r}^{pr}(\xi_{p})\colon X_{r}\longrightarrow X_{pr};\eta_{r}\mapsto u_{p,r}(\xi_{p}\otimes\eta_{r}).$
It follows that the family $\lambda(X):=\{\lambda_{p}(X_{p})\}_{p\in P}$ defines a concrete product system. Moreover, the family $\{\lambda_{p}\}_{p\in P}$ defines a unitary equivalence between $X$ and $\lambda(X)$. Hence $X$ and $\lambda(X)$ admit the same representations. In particular, if $\lambda^{\prime}$ is the Fock representation of $\lambda(X)$, then $\mathrm{C}^{*}(\lambda^{\prime})$ is unitarily equivalent to $\mathrm{C}^{*}(\lambda)$ by the unitary
$\bigoplus_{p\in P}\lambda_{p}\colon{\mathcal{F}}X\to{\mathcal{F}}\lambda(X);\xi_{p}\mapsto\lambda_{p}(\xi_{p}).$
Hence from an operator theoretic point of view, the concrete product systems that we will be using encompass Fowler’s product systems.
3. Fock covariant representations
3.1. Fock covariant representations
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Recall that $\lambda$ is injective and admits a normal coaction by $G$. Consider the induced ideal
${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}:=\left\langle\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}\right\rangle\lhd{\mathcal{T}}(X),$
and write ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ for the quotient of ${\mathcal{T}}(X)$ by ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$. Since ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$ is an induced ideal of ${\mathcal{T}}(X)$, then ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ inherits a coaction by $G$. The Fock covariant bundle of $X$ is the Fell bundle
${\mathcal{F}}{\mathcal{C}}X:=\Big{\{}[{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)]_{g}\Big{\}}_{g\in G},$
defined by the coaction by $G$ on ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$. A representation of $X$ that promotes to a representation of ${\mathcal{F}}{\mathcal{C}}X$ will be called a Fock covariant representation of $X$. We note that, the definition of the Fock covariant bundle does not depend on the choice of $G$. This follows by applying the first part of the proof of [47, Lemma 3.9] in this setting. Note also that since the bimodule properties are graded, every representation of ${\mathcal{F}}{\mathcal{C}}X$ is a representation of $X$.
Let $q_{\operatorname{c}}^{\operatorname{F}}\colon{\mathcal{T}}(X)\to{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ be the quotient map by the ideal ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$. Then
$q_{\operatorname{c}}^{\operatorname{F}}({\mathcal{P}}{\mathcal{S}}X)=\Big{\{}[{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)]_{g}\Big{\}}_{g\in G}={\mathcal{F}}{\mathcal{C}}X,$
and thus combining Proposition 2.2 with Proposition 2.7 we obtain a canonical *-isomorphism
${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)\simeq\mathrm{C}^{*}_{\max}({\mathcal{F}}{\mathcal{C}}X).$
By the definition of ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$ we have that
$\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{e}={\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{e}=\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}.$
By Proposition 2.1 we have that
$\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{g}=\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{g}\text{ for all }g\in G.$
Therefore $\lambda$ promotes to a representation of ${\mathcal{F}}{\mathcal{C}}X$ and we obtain the following commutative diagram
${{\mathcal{T}}(X)\simeq\mathrm{C}^{*}_{\max}({\mathcal{P}}{\mathcal{S}}X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda_{\ast}}$$\scriptstyle{q_{\operatorname{c}}^{\operatorname{F}}}$${{\mathcal{T}}_{\lambda}(X)}$${{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)\simeq\mathrm{C}^{*}_{\max}({\mathcal{F}}{\mathcal{C}}X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\dot{\lambda}}$
of canonical *-epimorphisms. Since $\lambda_{\ast}$ and $q_{\operatorname{c}}^{\operatorname{F}}$ are equivariant we have that so is $\dot{\lambda}$. On the other hand, since $\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{e}=\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}$, we have that $\dot{\lambda}$ is injective on the $e$-fiber. As the coaction by $G$ on $\lambda$ is normal, then so is the coaction by $G$ on $\dot{\lambda}$, and we conclude that
${\mathcal{T}}_{\lambda}(X)=\mathrm{C}^{*}(\dot{\lambda})\simeq\mathrm{C}^{*}_{\lambda}({\mathcal{F}}{\mathcal{C}}X).$
The Fock representation satisfies the following property. Let $p_{i},q_{i}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ such that $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e$ and $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P=\emptyset$. Proposition 2.4 yields
$\left(\lambda_{p_{1}}(\xi_{p_{1}})^{*}\right)^{\varepsilon}\lambda_{q_{1}}(\xi_{q_{1}})\cdots\lambda_{p_{n}}(\xi_{p_{n}})^{*}\lambda_{q_{n}}(\xi_{q_{n}})^{\varepsilon^{\prime}}X_{r}=(0)\text{ for all }r\in P,$
and therefore we have that
(3.1) $\mathbf{K}_{\emptyset,\lambda_{\ast}}=(0).$
We aim to give a characterisation for the equivariant Fock covariant injective representations of $X$. We begin with the following proposition.
Proposition 3.1.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\hat{t}$ be a representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$, and let $t$ be an equivariant representation of a product system $X$. Then $t$ is Fock covariant if and only if
$\ker\lambda_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\subseteq\ker t_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\text{ for every finite $\cap$-closed ${\mathcal{F}}\subseteq{\mathcal{J}}$}.$
Proof..
If $t$ is Fock covariant, then by definition $t_{\ast}$ factors through the canonical *-epimorphism $q_{\operatorname{c}}^{\operatorname{F}}\colon{\mathcal{T}}(X)\to{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$, and thus $\ker q_{\operatorname{c}}^{\operatorname{F}}\subseteq\ker t_{\ast}$. By the definition of ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$ we also have that $\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}=\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{e}$, and therefore
$\ker\lambda_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}=\ker q_{\operatorname{c}}^{\operatorname{F}}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\subseteq\ker t_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}},$
where we used that each $\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}$ is a C*-subalgebra of $[{\mathcal{T}}(X)]_{e}$.
For the converse, recall that the fixed point algebra $[{\mathcal{T}}(X)]_{e}$ is the inductive limit of $\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}$ for $\cap$-closed families ${\mathcal{F}}\subseteq{\mathcal{J}}$. By the properties of the inductive limits we have that
$\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{e}$ $=\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}=\overline{\bigcup_{{\mathcal{F}}}\left(\ker\lambda_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\right)}$
$\subseteq\overline{\bigcup_{{\mathcal{F}}}\left(\ker t_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\right)}=\ker t_{\ast}\cap[{\mathcal{T}}(X)]_{e},$
where we used the assumption for the inclusion. By combining Proposition 2.1 with Proposition 2.7 we get that $t$ induces a representation of ${\mathcal{F}}{\mathcal{C}}X$, i.e., $t$ is Fock covariant, and the proof is complete. ∎
We can now provide the characterisation of equivariant Fock covariant injective representations of $X$.
Theorem 3.2.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\hat{t}$ be a representation of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. An equivariant injective representation $t$ of $X$ is Fock covariant if and only if $t$ satisfies the following conditions:
  1. (i)
    $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$.
  2. (ii)
    For any $\cap$-closed =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}\subseteq{\mathcal{J}}$ such that $\bigcup_{i=1}^{n}\mathbf{x}_{i}\neq\emptyset$, and any $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$, with $i=1,\dots,n$, the following property holds:
    if $\sum\limits_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(X_{r})=(0)$ for all $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then $\sum\limits_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=0$.
Proof..
First we note that the Fock representation $\lambda$ satisfies conditions (i) and (ii). Condition (i) for $\lambda$ is shown in (3.1). For condition (ii), let =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ be a $\cap$-closed finite subset of ${\mathcal{J}}$ such that $\bigcup_{i=1}^{n}\mathbf{x}_{i}\neq\emptyset$, and let $b_{\mathbf{x}_{i}}$ be in $\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ such that
$\sum_{i:r\in\mathbf{x}_{i}}\lambda_{\ast}(b_{\mathbf{x}_{i}})\lambda_{r}(\xi_{r})=0\text{ for all }r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}.$
Recall that $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ by Proposition 2.6. Therefore, if $r\notin\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. On the other hand, if $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then we have that
$\lambda_{r}\left(\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}\right)=\sum_{i:r\in\mathbf{x}_{i}}\lambda_{r}(\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r})=\sum_{i:r\in\mathbf{x}_{i}}\lambda_{\ast}(b_{\mathbf{x}_{i}})\lambda_{r}(\xi_{r})=0,$
where in the first equality we used that $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ from Proposition 2.6, and in the second equality we used Proposition 2.5. By injectivity of $\lambda_{r}$, we have that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. This concludes that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})=0$, as required.
Now suppose that $t$ is Fock covariant. Recall that
$\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{g}=\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{g}\text{ for all }g\in G,$
by the definition of ${\mathcal{J}}_{\operatorname{c}}^{\operatorname{F}}$. Since $t$ is Fock covariant, we have that $t_{\ast}$ factors through the quotient map $q_{\operatorname{c}}^{\operatorname{F}}\colon{\mathcal{T}}(X)\to{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$, and therefore
$\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{g}=\ker q_{\operatorname{c}}^{\operatorname{F}}\cap[{\mathcal{T}}(X)]_{g}\subseteq\ker t_{\ast}\cap[{\mathcal{T}}(X)]_{g}\text{ for all }g\in G.$
For condition (i), we have that
$\mathbf{K}_{\emptyset,\hat{t}_{\ast}}\subseteq\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}\subseteq\ker t_{\ast}\cap[{\mathcal{T}}(X)]_{e},$
where we used that $\lambda$ satisfies condition (i) in the first inclusion. It thus follows that
$\mathbf{K}_{\emptyset,t_{\ast}}=t_{\ast}(\mathbf{K}_{\emptyset,\hat{t}_{\ast}})=(0),$
as required.
For condition (ii), let =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ be a $\cap$-closed finite subset of ${\mathcal{J}}$ such that $\bigcup_{i=1}^{n}\mathbf{x}_{i}\neq\emptyset$, and let $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ such that
$\sum_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(X_{r})=(0)\text{ for all }r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}.$
Fix $r\in P$. If $r\not\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, then Proposition 2.6 yields
$\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})X_{r}=(0).$
On the other hand, if $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ then for every $\xi_{r}\in X_{r}$ we have that
$t_{r}\left(\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}\right)=\sum_{i:r\in\mathbf{x}_{i}}t_{r}(\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r})=\sum_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(\xi_{r})=0,$
where in the first equality we used that $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})\xi_{r}=0$ whenever $r\notin{\mathbf{x}}_{i}$ from Proposition 2.6, and in the second equality we used Proposition 2.5. Injectivity of $t$ implies that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$. We conclude that $\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})=0$ and thus
$\sum_{i=1}^{n}b_{\mathbf{x}_{i}}\in\ker\lambda_{\ast}\cap[{\mathcal{T}}(X)]_{e}\subseteq\ker t_{\ast}\cap[{\mathcal{T}}(X)]_{e}.$
Therefore $\sum_{i=1}^{n}t_{\ast}(b_{{\mathbf{x}}_{i}})=0$, as required. This completes the proof of the one direction.
For the converse, suppose that $t$ satisfies conditions (i) and (ii). By Proposition 3.1, it suffices to prove that $\ker\lambda_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}\subseteq\ker t_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}$ for every finite $\cap$-closed ${\mathcal{F}}\subseteq{\mathcal{J}}$. If ${}{\mathcal{F}}=\{\emptyset\}$, then condition (i) for $\lambda$ and $t$ yields
$\ker\lambda_{\ast}\cap\mathbf{B}_{\{\emptyset\},\hat{t}_{\ast}}=\ker\lambda_{\ast}\cap\mathbf{K}_{\emptyset,\hat{t}_{*}}=\mathbf{K}_{\emptyset,\hat{t}_{*}}=\ker t_{\ast}\cap\mathbf{K}_{\emptyset,\hat{t}_{*}}=\ker t_{\ast}\cap\mathbf{B}_{\{\emptyset\},\hat{t}_{\ast}},$
where we used that $\mathbf{K}_{\emptyset,\hat{t}_{*}}=\mathbf{B}_{\{\emptyset\},\hat{t}_{\ast}}$ is a C*-subalgebra of $[{\mathcal{T}}(X)]_{e}$. Next suppose that =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ such that $\bigcup_{i=1}^{n}{\mathbf{x}}_{i}\neq\emptyset$, and let $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$ such that
$\sum_{i=1}^{n}b_{\mathbf{x}_{i}}\in\ker\lambda_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}}.$
For $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ we have that
$\sum_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(\xi_{r})=t_{r}\left(\sum_{i:r\in\mathbf{x}_{i}}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}\right)=t_{r}\left(\sum_{i=1}^{n}\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}\right)=0,$
where in the first equality we used Proposition 2.5, and on the second equality we used that $\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r}=0$ if $r\notin\mathbf{x}_{i}$ from Proposition 2.6. Condition (iii) now yields that $\sum_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=0$, and thus
$\sum_{i=1}^{n}b_{\mathbf{x}_{i}}\in\ker t_{\ast}\cap\mathbf{B}_{{\mathcal{F}},\hat{t}_{\ast}},$
as required. ∎
Remark 3.3.
We note that, if $t$ is an equivariant Fock covariant injective representation of $X$, then the induced *-representation ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)\to\mathrm{C}^{*}(t)$ is injective on every $\mathbf{K}_{\bullet}$-core of ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$. Indeed, let $b_{{\mathbf{x}}}\in\mathbf{K}_{{\mathbf{x}},\hat{t}}$ for some $\emptyset\neq{\mathbf{x}}\in{\mathcal{J}}$ such that $\lambda_{\ast}(b_{{\mathbf{x}}})\not=0$, where $\mathrm{C}^{*}(\hat{t})={\mathcal{T}}(X)$. Then there exists a $\xi_{r}\in X_{r}$ such that $\lambda_{\ast}(b_{{\mathbf{x}}})\xi_{r}\neq 0$. By Proposition 2.5 we deduce that
$t_{\ast}(b_{{\mathbf{x}}})t_{r}(\xi_{r})=t_{r}(\lambda_{\ast}(b_{{\mathbf{x}}})\xi_{r})\neq 0,$
as $t$ is injective, and therefore $t_{\ast}(b_{{\mathbf{x}}})\neq 0$.
3.2. Compactly aligned product systems over right LCM semigroups
A unital semigroup $P$ is said to be a right LCM semigroup if it is left cancellative and satisfies Clifford’s condition, i.e., for every $p,q\in P$ with $pP\cap qP\neq\emptyset$, there exists a $w\in P$ such that $pP\cap qP=wP$. The element $w$ is referred to as a right least common multiple or right LCM of $p$ and $q$. If an element $w\in P$ is a right least common multiple of $p,q\in P$, then so is $wr$ for every $r\in P^{*}:=P\cap P^{-1}$. It follows that $P$ is a right LCM semigroup if and only if ${\mathcal{J}}=\{pP\mid p\in P\}\bigcup\{\emptyset\}$.
Right LCM semigroups include as a special case the quasi-lattice ordered semigroups considered in [42]. In particular, a unital right LCM semigroup is a quasi-lattice if and only if $P^{*}=\{e\}$, and in this case a right common least multiple is unique. Artin monoids [6], the Baumslag-Solitar monoids $B(m,n)^{+}$ [21, 38, 49], and the semigroup $R\rtimes R^{\times}$ of affine transformations of an integral domain $R$ that satisfies the GCD condition [37, 43] are examples of right LCM semigroups.
Fowler’s original work related to product systems over quasi-lattices [18]. Product systems over right LCM semigroups were introduced and studied by Kwaśniewski and Larsen [32, 33], extending the construction of Fowler, and they have been investigated further in [11, 23]. The interest lies in that they retain several of the structural properties from the single C*-correspondence case. With the addition of one further property, the generated C*-algebras admit a Wick ordering. A product system $X$ over $P$ in the sense of Fowler gives rise to *-homomorphisms
$i_{p}^{pq}\colon{\mathcal{L}}(X_{p})\to{\mathcal{L}}(X_{pq})\text{ where }i_{p}^{pq}(S):=u_{p,q}(S\otimes{\operatorname{id}}_{X_{q}})u_{p,q}^{*}\text{ for all }S\in{\mathcal{L}}(X_{p}).$
We say that $X$ is compactly aligned if $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\in{\mathcal{K}}(X_{w})$ for all $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$, whenever $pP\cap qP=wP$. It was established in [11] that this definition is independent of the choice of $w$. For a representation $t$ of a product system $X$ we use the notation
$t^{(p)}\colon{\mathcal{K}}(X_{p})\to\mathrm{C}^{*}(t);\theta^{X_{p}}_{\xi_{p},\eta_{p}}\mapsto t_{p}(\xi_{p})t_{p}(\eta_{p})^{*},$
for the induced *-representation of the compact operators ${\mathcal{K}}(X_{p})$.
Henceforth let us assume that the right LCM semigroup $P$ is in a group $G$.
Proposition 3.4.
Let $P$ be a unital right LCM semigroup that is a subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$ in the sense of Fowler. Then $X$ is compactly aligned if and only if for every $p,q\in P$ with $pP\cap qP=wP$ we have that
$\lambda^{(p)}({\mathcal{K}}(X_{p}))\lambda^{(q)}({\mathcal{K}}(X_{q}))\subseteq\lambda^{(w)}({\mathcal{K}}(X_{w})).$
Proof..
First suppose that $X$ is compactly aligned, and let $p,q\in P$ with $pP\cap qP=wP$. Let $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$. By compact alignment we have that $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\in{\mathcal{K}}(X_{w})$. Hence for $r\in wP$ we get
$\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})\xi_{r}$ $=\lambda^{(p)}(k_{p})(i_{q}^{r}(k_{q})\xi_{r})=i_{p}^{r}(k_{p})(i_{q}^{r}(k_{q})\xi_{r})$
$=i_{w}^{r}(i_{p}^{w}(k_{p}))i_{w}^{r}(i_{q}^{w}(k_{q}))\xi_{r}=i_{w}^{r}(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q}))\xi_{r}=\lambda^{(w)}(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q}))\xi_{r}.$
On the other hand for $r\not\in wP$, we have that $\lambda^{(w)}(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q}))\xi_{r}=0$. If $r\notin qP$ then $\lambda^{(q)}(k_{q})\xi_{r}=0$. If $r\in qP$ then $\lambda^{(q)}(k_{q})\xi_{r}\in X_{r}$, but then it has to be that $r\notin pP$, and so $\lambda^{(p)}(k_{p})(\lambda^{(q)}(k_{q})\xi_{r})=0$. In all cases, for $r\notin wP$ we have that
$\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})\xi_{r}=0=\lambda^{(w)}(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q}))\xi_{r}.$
Therefore, we conclude that
$\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})=\lambda^{(w)}(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q}))\in\lambda^{(w)}({\mathcal{K}}(X_{w})).$
For the converse, let $p,q\in P$ such that $pP\cap qP=wP$ for some $w\in P$, and let $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$. Since
$\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})\in\lambda^{(w)}({\mathcal{K}}(X_{w})),$
there exists a $k_{w}\in{\mathcal{K}}(X_{w})$ such that $\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})=\lambda^{(w)}(k_{w})$. For every $\xi_{w}\in X_{w}$ we get that
$i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\xi_{w}=\lambda^{(p)}(k_{p})\lambda^{(q)}(k_{q})\xi_{w}=\lambda^{(w)}(k_{w})\xi_{w}=k_{w}\xi_{w}.$
Therefore we have that $i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})=k_{w}\in{\mathcal{K}}(X_{w})$, and the proof is complete. ∎
Let $P$ be a unital right LCM semigroup that is a subsemigroup of a discrete group $G$ and let $X$ be a compactly aligned product system over $P$ in the sense of Fowler. Let $t$ be a representation of $X$. We say that $t$ is Nica covariant if and only if
$t^{(p)}(k_{p})t^{(q)}(k_{q})=\begin{cases}t^{(w)}\left(i_{p}^{w}(k_{p})i_{q}^{w}(k_{q})\right)&\text{if }pP\cap qP=wP,\\ 0&\text{if }pP\cap qP=\emptyset,\end{cases}$
for all $k_{p}\in{\mathcal{K}}(X_{p})$ and $k_{q}\in{\mathcal{K}}(X_{q})$. Because of linearity and continuity of $t^{(p)},t^{(q)}$ and $t^{(w)}$, we have that $t$ is Nica covariant if and only if
$t^{(p)}\left(\theta^{X_{p}}_{\xi_{p},\eta_{p}}\right)t^{(q)}\left(\theta^{X_{q}}_{\xi_{q},\eta_{q}}\right)=\begin{cases}t^{(w)}\left(i_{p}^{w}\left(\theta^{X_{p}}_{\xi_{p},\eta_{p}}\right)i_{q}^{w}\left(\theta^{X_{q}}_{\xi_{q},\eta_{q}}\right)\right)&\text{if }pP\cap qP=wP,\\ 0&\text{if }pP\cap qP=\emptyset,\end{cases}$
for all $\xi_{p},\eta_{p}\in X_{p}$ and $\xi_{q},\eta_{q}\in X_{q}$. We write ${\mathcal{N}}{\mathcal{T}}(X)$ for the universal C*-algebra with respect to the Nica covariant representations of $X$.
In [11, Proposition 2.4] it is shown that, if $w\in P$ and $r\in P^{*}$, then $i_{w}^{wr}(k_{w})\in{\mathcal{K}}(X_{wr})$ and $t^{(wr)}(i_{w}^{wr}(k_{w}))=t^{(w)}(k_{w})\text{ for all }k_{w}\in{\mathcal{K}}(X_{w})$, when $t$ is Nica covariant. Moreover, in the discussion following [11, Definition 2.9] it is shown that
(3.2) $t_{p}(X_{p})^{*}t_{q}(X_{q})=(0)\text{ for }p,q\in P\text{ such that }pP\cap qP=\emptyset,$
and that
(3.3) $t_{p}(X_{p})^{*}t_{q}(X_{q})\subseteq[t_{r}(X_{r})t_{s}(X_{s})^{*}]\text{ for }wP=pP\cap qP\text{ and }r=p^{-1}w,s=q^{-1}w.$
Consequently, Nica covariance does not depend on the choice of a right least common multiple and also $\mathrm{C}^{*}(t)$ admits a Wick ordering in the sense that
$\mathrm{C}^{*}(t)=\overline{\operatorname{span}}\{t_{p}(X_{p})t_{q}(X_{q})^{*}\mid p,q\in P\}.$
Moreover we have that $\mathbf{K}_{pP,t_{\ast}}=t^{(p)}({\mathcal{K}}(X_{p}))$. Therefore, if =${\mathcal{F}}=\{p_{1}P,\dots,p_{n}P\}$ is a finite $\cap$-closed subset of ${\mathcal{J}}$, then we have that
$\mathbf{B}_{{\mathcal{F}},t_{\ast}}=\sum_{i=1}^{n}t^{(p_{i})}({\mathcal{K}}(X_{p_{i}})),$
irrespectively of the choice of the elements $p_{1},\dots,p_{n}$. It is implicit in [11], and proven in a more general context in [23, Theorem 6.4], that the cores $\mathbf{K}_{p_{i}P,\lambda_{\ast}}$ for distinct $p_{i}P$’s are linearly independent in $\mathbf{B}_{{\mathcal{F}},\lambda_{\ast}}$.
It is remarked in [11] that ${\mathcal{N}}{\mathcal{T}}(X)$ admits a coaction since Nica covariance is a graded relation. Let ${\mathcal{N}}X$ be the induced Fell bundle in ${\mathcal{N}}{\mathcal{T}}(X)$. By [11, Proposition 4.3] we have that
${\mathcal{N}}{\mathcal{T}}(X)\simeq\mathrm{C}^{*}_{\max}({\mathcal{N}}X)\quad\text{and}\quad{\mathcal{T}}_{\lambda}(X)\simeq\mathrm{C}^{*}_{\lambda}({\mathcal{N}}X),$
by canonical *-isomorphisms. In particular the second *-isomorphism induces a Fell bundle isomorphism between ${\mathcal{N}}X$ and ${\mathcal{F}}{\mathcal{C}}X$. This shows that a representation of $X$ is Fock covariant if and only if it is Nica covariant. Below we will give an alternative proof that squares with our characterisation of Fock covariance. We will need the following proposition.
Proposition 3.5.
Let $P$ be a unital right LCM semigroup that is a subsemigroup of a discrete group $G$ and let $X$ be a compactly aligned product system over $P$ in the sense of Fowler. If $t$ is a Nica covariant representation of $X$, then
$\left(t_{p_{1}}(X_{p_{1}})^{*}\right)^{\varepsilon}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})^{\varepsilon^{\prime}}=(0)$
for $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ such that $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P=\emptyset$. In particular, we have that $\mathbf{K}_{\emptyset,t_{\ast}}=(0)$.
Proof..
We first consider the case $\varepsilon=\varepsilon^{\prime}=1$. In this case we have that $q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P=\emptyset$ and we wish to show that
$t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})=(0).$
We proceed in steps. If $p_{n}P\cap q_{n}P=\emptyset$, then Nica covariance of $t$ implies that
$t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})=(0)$
by equation (3.2), which gives the desired result. If $p_{n}P\cap q_{n}P\not=\emptyset$, then choose $w_{1}\in P$ with $p_{n}P\cap q_{n}P=w_{1}P$, and (3.3) implies that
$t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})\subseteq[t_{r_{1}}(X_{r_{1}})t_{s_{1}}(X_{s_{1}})^{*}]\quad\text{for}\quad r_{1}=p_{n}^{-1}w_{1},s_{1}=q_{n}^{-1}w_{1}.$
Hence we obtain the inclusion
$t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})\subseteq$
$\hskip 85.35826pt\subseteq[t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n-1}}(X_{p_{n-1}})^{*}t_{q_{n-1}}(X_{q_{n-1}})t_{r_{1}}(X_{r_{1}})t_{s_{1}}(X_{s_{1}})^{*}]$
$\hskip 85.35826pt\subseteq[t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n-1}}(X_{p_{n-1}})^{*}t_{q_{n-1}r_{1}}(X_{q_{n-1}r_{1}})t_{s_{1}}(X_{s_{1}})^{*}].$
If $p_{n-1}P\cap q_{n-1}r_{1}P=\emptyset$, then Nica covariance gives the desired result as
$t_{p_{n-1}}(X_{p_{n-1}})^{*}t_{q_{n-1}r_{1}}(X_{q_{n-1}r_{1}})=(0).$
If $p_{n-1}P\cap q_{n-1}r_{1}P\not=\emptyset$, then choose $w_{2}\in P$ such that $p_{n-1}P\cap q_{n-1}r_{1}P=w_{2}P$, and (3.3) implies that
$t_{p_{n-1}}(X_{p_{n-1}})^{*}t_{q_{n-1}r_{1}}(X_{q_{n-1}r_{1}})\subseteq[t_{r_{2}}(X_{r_{2}})t_{s_{2}}(X_{s_{2}})^{*}]\quad\text{for}\quad r_{2}=p_{n-1}^{-1}w_{2},s_{2}=r_{1}^{-1}q_{n-1}^{-1}w_{2}.$
Hence we obtain the inclusion
$t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})\subseteq$
$\hskip 85.35826pt\subseteq[t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n-1}}(X_{p_{n-1}})^{*}t_{q_{n-1}r_{1}}(X_{q_{n-1}r_{1}})t_{s_{1}}(X_{s_{1}})^{*}]$
$\hskip 85.35826pt\subseteq[t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{q_{n-2}}(X_{q_{n-2}})t_{r_{2}}(X_{r_{2}})t_{s_{2}}(X_{s_{2}})^{*}t_{s_{1}}(X_{s_{1}})^{*}]$
$\hskip 85.35826pt\subseteq[t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n-2}}(X_{p_{n-2}})^{*}t_{q_{n-2}r_{2}}(X_{q_{n-2}r_{2}})t_{s_{1}s_{2}}(X_{s_{1}s_{2}})^{*}].$
Continuing inductively we either get a zero space at a step, or we have obtained elements $w_{k},r_{k},s_{k}\in P$, with $k=2,\dots,n-1$, such that
$w_{k}P$ $=p_{n-(k-1)}P\cap q_{n-(k-1)}r_{k-1}P,\,r_{k}=p_{n-(k-1)}^{-1}w_{k},\,s_{k}=r_{k-1}^{-1}q_{n-(k-1)}^{-1}w_{k},$
and the inclusion
$t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})\subseteq[t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}r_{n-1}}(X_{q_{1}r_{n-1}})t_{s_{1}\cdots s_{n-1}}(X_{s_{1}\cdots s_{n-1}})^{*}].$
In the latter case we claim that
$p_{1}P\cap q_{1}r_{n-1}P=\emptyset,$
and so Nica covariance will imply the required identity. In order to reach a contradiction suppose that $p_{1}P\cap q_{1}r_{n-1}P\neq\emptyset$, so there exists a $q\in P$ such that $p:=r_{n-1}^{-1}q_{1}^{-1}p_{1}q\in P$. By the construction of the $r_{k}$ and the $s_{k}$ for $k=1,2,\dots,n-1$ we have that
$s_{1}$ $=q_{n}^{-1}w_{1}=q_{n}^{-1}p_{n}r_{1}$
$s_{2}$ $=r_{1}^{-1}q_{n-1}^{-1}w_{2}=r_{1}^{-1}q_{n-1}^{-1}p_{n-1}r_{2}$
$\vdots$
$s_{k}$ $=r_{k-1}^{-1}q_{n-(k-1)}^{-1}w_{k}=r_{k-1}^{-1}q_{n-(k-1)}^{-1}p_{n-(k-1)}r_{k}$
$\vdots$
$s_{n-1}$ $=r_{n-2}^{-1}q_{2}^{-1}w_{n-1}=r_{n-2}^{-1}q_{2}^{-1}p_{2}r_{n-1}.$
Therefore for every $k=1,2\dots,n-1$ we have that
$s_{1}s_{2}\cdots s_{k}\cdots s_{n-1}p$ $=q_{n}^{-1}p_{n}q_{n-1}^{-1}p_{n-1}\cdots q_{n-(k-1)}^{-1}p_{n-(k-1)}\cdot\left(r_{k}s_{k+1}\cdots s_{n-1}p\right)$
$\in q_{n}^{-1}\cdot p_{n}\cdots q_{n-(k-1)}^{-1}\cdot p_{n-(k-1)}\cdot P.$
Because of the choice of $p=r_{n-1}^{-1}q_{1}^{-1}p_{1}q$ we also have that
$s_{1}\cdots s_{n-2}s_{n-1}p$ $=q_{n}^{-1}p_{n}\cdots q_{2}^{-1}p_{2}r_{n-1}\cdot\left(r_{n-1}^{-1}q_{1}^{-1}p_{1}q\right)\in q_{n}^{-1}\cdot p_{n}\cdots q_{1}^{-1}\cdot p_{1}\cdot P.$
Therefore we have that $s_{1}\cdots s_{n-2}s_{n-1}p\in q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P=\emptyset$ which is a contradiction, and the proof for the first case is complete.
Suppose now that $\varepsilon=1$ and $\varepsilon^{\prime}=0$. We have that $p_{n}q_{n-1}^{-1}\dots q_{1}^{-1}p_{1}P=\emptyset$ and we wish to show that
$t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{q_{n-1}}(X_{q_{n-1}})t_{p_{n}}(X_{p_{n}})^{*}=(0).$
Note that we must also have that $q_{n-1}^{-1}p_{n-1}\dots q_{1}^{-1}p_{1}P=\emptyset$, and therefore the previous case implies that
$t_{p_{1}}(X_{p_{1}})^{*}t_{q_{1}}(X_{q_{1}})\cdots t_{p_{n-1}}(X_{p_{n-1}})^{*}t_{q_{n-1}}(X_{q_{n-1}})=(0),$
which completes the proof of this case.
Suppose now that $\varepsilon=0$ and $\varepsilon^{\prime}=1$. We have that $q_{n}^{-1}p_{n}\dots p_{2}q_{1}^{-1}P=\emptyset$ and we wish to show that
$t_{q_{1}}(X_{q_{1}})t_{p_{2}}(X_{p_{2}})^{*}\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})=(0).$
Since $q_{n}^{-1}p_{n}\dots q_{2}^{-1}p_{2}P=q_{n}^{-1}p_{n}\dots p_{2}q_{1}^{-1}P=\emptyset$, the first case implies that
$t_{p_{2}}(X_{p_{2}})^{*}t_{q_{2}}(X_{q_{2}})\cdots t_{p_{n}}(X_{p_{n}})^{*}t_{q_{n}}(X_{q_{n}})=(0),$
which completes the proof of this case.
Finally, suppose that $\varepsilon=\varepsilon^{\prime}=0$. We have that $p_{n}q_{n-1}^{-1}\dots p_{2}q_{1}^{-1}P=\emptyset$ and we wish to show that
$t_{q_{1}}(X_{q_{1}})t_{p_{2}}(X_{p_{2}})^{*}\cdots t_{q_{n-1}}(X_{q_{n-1}})t_{p_{n}}(X_{p_{n}})^{*}=(0).$
Since $p_{n}q_{n-1}^{-1}\dots q_{2}^{-1}p_{2}P=p_{n}q_{n-1}^{-1}\dots p_{2}q_{1}^{-1}P=\emptyset$, the second case implies that
$t_{p_{2}}(X_{p_{2}})^{*}t_{q_{2}}(X_{q_{2}})\cdots t_{q_{n-1}}(X_{q_{n-1}})t_{p_{n}}(X_{p_{n}})^{*}=(0),$
and the proof is complete. ∎
We now pass to the connection between Fock covariance and Nica covariance through our characterisation.
Proposition 3.6.
Let $P$ be a unital right LCM semigroup that is a subsemigroup of a discrete group $G$ and let $X$ be a compactly aligned product system over $P$ in the sense of Fowler. An equivariant injective representation of $X$ is Fock covariant if and only if it is Nica covariant.
Proof..
Fix the representation $\hat{t}$ of $X$ such that ${\mathcal{T}}(X)=\mathrm{C}^{*}(\hat{t})$. First suppose that $t$ is an equivariant Fock covariant injective representation of $X$, so that it satisfies conditions (i) and (ii) of Theorem 3.2. Fix $p,q\in P$. We will show that
$t_{p}(\xi_{p})t_{p}(\eta_{p})^{*}t_{q}(\xi_{q})t_{q}(\eta_{q})^{*}=\begin{cases}t^{(w)}\left(i_{p}^{w}\left(\theta^{X_{p}}_{\xi_{p},\eta_{p}}\right)i_{q}^{w}\left(\theta^{X_{q}}_{\xi_{q},\eta_{q}}\right)\right)&\text{if }pP\cap qP=wP,\\ 0&\text{if }pP\cap qP=\emptyset,\end{cases}$
for all $\xi_{p},\eta_{p}\in X_{p}$ and $\xi_{q},\eta_{q}\in X_{q}$. Towards this end, set
$\mathbf{x}:=qq^{-1}pp^{-1}P=pP\cap qP.$
If $\mathbf{x}=\emptyset$, then condition (i) of Theorem 3.2 implies that
$t_{p}(\xi_{p})t_{p}(\eta_{p})^{*}t_{q}(\xi_{q})t_{q}(\eta_{q})^{*}\in\mathbf{K}_{\emptyset,t_{\ast}}=(0),$
as required. On the other hand, if $\mathbf{x}=wP$ for some $w\in P$ set
$b_{{\mathbf{x}}}:=\hat{t}_{p}(\xi_{p})\hat{t}_{p}(\eta_{p})^{*}\hat{t}_{q}(\xi_{q})\hat{t}_{q}(\eta_{q})^{*}\quad\text{and}\quad k_{w}:=i_{p}^{w}\left(\theta^{X_{p}}_{\xi_{p},\eta_{p}}\right)i_{q}^{w}\left(\theta^{X_{q}}_{\xi_{q},\eta_{q}}\right)\in{\mathcal{K}}(X_{w}).$
We have to show that $t_{\ast}(b_{{\mathbf{x}}})=t^{(w)}\left(k_{w}\right)$. We will use that $t^{(w)}(k_{w})=t_{\ast}(\hat{t}^{(w)}(k_{w}))$. Since ${\mathbf{x}}=qq^{-1}pp^{-1}P$ we have that $b_{{\mathbf{x}}}\in\mathbf{K}_{{\mathbf{x}},\hat{t}_{\ast}}$. Moreover, since ${\mathbf{x}}=wP=ww^{-1}P$ we have that
$\hat{t}^{(w)}(k_{w})\in[\hat{t}_{w}(X_{w})\hat{t}_{w}(X_{w})^{*}]\subseteq\mathbf{K}_{\mathbf{x},\hat{t}_{\ast}},$
and therefore
$t_{\ast}(b_{{\mathbf{x}}})-t_{\ast}(\hat{t}^{(w)}(k_{w}))\in\mathbf{K}_{{\mathbf{x}},t_{\ast}}.$
Fix $r\in\mathbf{x}$. Then we have that
$\lambda_{\ast}(b_{{\mathbf{x}}})\xi_{r}$ $=\lambda^{(p)}\left(\theta^{X_{p}}_{\xi_{p},\eta_{p}}\right)\lambda^{(q)}\left(\theta^{X_{q}}_{\xi_{q},\eta_{q}}\right)\xi_{r}=i_{p}^{r}\left(\theta^{X_{p}}_{\xi_{p},\eta_{p}}\right)i_{q}^{r}\left(\theta^{X_{q}}_{\xi_{q},\eta_{q}}\right)\xi_{r}=i_{w}^{r}\left(k_{w}\right)\xi_{r},$
and thus by Proposition 2.5 we have
$t_{\ast}(b_{{\mathbf{x}}})t_{r}(\xi_{r})=t_{r}(\lambda_{\ast}(b_{{\mathbf{x}}})\xi_{r})=t_{r}\left(i_{w}^{r}\left(k_{w}\right)\xi_{r}\right).$
On the other hand, by Proposition 2.5 we have that
$t^{(w)}(k_{w})t_{r}(\xi_{r})=t_{\ast}(\hat{t}^{(w)}(k_{w}))t_{r}(\xi_{r})=t_{r}(\lambda_{\ast}(\hat{t}^{(w)}(k_{w}))\xi_{r})=t_{r}(\lambda^{(w)}(k_{w})\xi_{r})=t_{r}(i_{w}^{r}(k_{w})\xi_{r}).$
Therefore we conclude that
$\left(t_{\ast}(b_{{\mathbf{x}}})-t_{\ast}(\hat{t}^{(w)}(k_{w}))\right)t_{r}(\xi_{r})=\left(t_{\ast}(b_{{\mathbf{x}}})-t^{(w)}(k_{w})\right)t_{r}(\xi_{r})=0\text{ for all }r\in{\mathbf{x}}.$
Condition (ii) for $t$ then implies that $t_{\ast}(b_{{\mathbf{x}}})=t_{\ast}(\hat{t}^{(w)}(k_{w}))=t^{(w)}(k_{w})$, as required.
For the converse, suppose that $t$ is an equivariant Nica covariant injective representation of $X$, and we will show that $t$ satisfies conditions (i) and (ii) of Theorem 3.2. The fact that $t$ satisfies condition (i) is already verified in Proposition 3.5. We will show that $t$ satisfies condition (ii). In order to fix notation let the commutative diagrams of canonical *-epimorphisms
[…]
induced by $\lambda$ and $t$, and let $\tilde{t}$ be the representation of $X$ such that ${\mathcal{N}}{\mathcal{T}}(X)=\mathrm{C}^{*}(\tilde{t})$. We will use the following note for an element $b\in\mathbf{K}_{pP,\hat{t}_{\ast}}$. Since
$q_{\operatorname{c}}^{\rm{N}}(b)\in q_{\operatorname{c}}^{\rm N}(\mathbf{K}_{pP,\hat{t}_{\ast}})=\mathbf{K}_{pP,\tilde{t}_{\ast}}=\tilde{t}^{(p)}({\mathcal{K}}(X_{p})),$
there exists a $k_{p}\in{\mathcal{K}}(X_{p})$ such that $q_{\operatorname{c}}^{\rm{N}}(b)=\tilde{t}^{(p)}(k_{p})$. In particular we have that
$t_{\ast}(b)=\dot{t}\circ q_{\operatorname{c}}^{\rm{N}}(b)=\dot{t}(\tilde{t}^{(p)}(k_{p}))=t^{(p)}(k_{p}).$
Likewise we have that
$\lambda_{\ast}(b)=\dot{\lambda}\circ q_{\operatorname{c}}^{\rm{N}}(b)=\dot{\lambda}(\tilde{t}^{(p)}(k_{p}))=\lambda^{(p)}(k_{p}).$
For condition (ii), first let =${\mathcal{F}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ be a $\cap$-closed subset of ${\mathcal{J}}$ where $\mathbf{x}_{i}=p_{i}P$ and $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$. Note that without loss of generality we may suppose that ${\mathbf{x}}_{i}\not={\mathbf{x}}_{j}$ for $i\not=j$ and hence the C*-subalgebras $\mathbf{K}_{p_{i}P,\lambda_{\ast}}$’s are linearly independent in $\mathbf{B}_{{\mathcal{F}},\lambda_{\ast}}$, by [23, Theorem 6.4]. Suppose that
$\sum_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(X_{r})=(0)\text{ for all }r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}.$
We will show that $q_{\operatorname{c}}^{\rm{N}}(b_{\mathbf{x}_{i}})=0$ for all $i=1,\dots,n$, and thus
$\sum_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=\sum_{i=1}^{n}\dot{t}\left(q_{\operatorname{c}}^{\rm{N}}(b_{\mathbf{x}_{i}})\right)=0.$
From the comments above, for each $i=1,\dots,n$ we may pick a $k_{p_{i}}\in{\mathcal{K}}(X_{p_{i}})$ such that $q_{\operatorname{c}}^{\rm{N}}(b_{\mathbf{x}_{i}})=\tilde{t}^{(p_{i})}(k_{p_{i}})$. Then $\lambda_{\ast}(b_{{\mathbf{x}}_{i}})=\lambda^{(p_{i})}(k_{p_{i}})$ for all $i=1,\dots,n$, and thus for each $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ we have that
$t_{r}\left(\sum_{i=1}^{n}\lambda^{(p_{i})}(k_{p_{i}})\xi_{r}\right)=\sum_{i:r\in\mathbf{x}_{i}}t_{r}(\lambda^{(p_{i})}(k_{p_{i}})\xi_{r})=\sum_{i:r\in\mathbf{x}_{i}}t_{r}(\lambda_{\ast}(b_{\mathbf{x}_{i}})\xi_{r})=\sum_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(\xi_{r})=0,$
by using Proposition 2.5. Injectivity of $t_{r}$ now yields that $\sum_{i=1}^{n}\lambda^{(p_{i})}(k_{p_{i}})\xi_{r}=0$ for all $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$. On the other hand we have that $\lambda^{(p_{i})}(k_{p_{i}})\xi_{r}=0$ for all $r\not\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ and $i=1,\dots,n$, and therefore we obtain that
$\sum_{i=1}^{n}\lambda^{(p_{i})}(k_{p_{i}})=0.$
As the $\mathbf{K}_{\bullet}$-cores are linearly independent by [23, Theorem 6.4], we deduce that $\lambda^{(p_{i})}(k_{p_{i}})=0$ for all $i=1,\dots,n$, and thus $k_{p_{i}}=0$ for all $i=1,\dots,n$, from the injectivity of $\lambda^{(p_{i})}$. In particular we have that
$q_{\operatorname{c}}^{\rm{N}}(b_{\mathbf{x}_{i}})=\tilde{t}^{(p_{i})}(k_{p_{i}})=0\text{ for all }i=1,\dots,n,$
as required. To finish the proof, next let =${\mathcal{F}}=\{{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}\}\cup\{\emptyset\}$ such that ${\mathbf{x}}_{i}\not={\mathbf{x}}_{j}$ for $i\not=j$, and ${\mathbf{x}}_{i}=p_{i}P$ for some $p_{i}\in P$. Let $b_{\emptyset},b_{{\mathbf{x}}_{1}},\dots,b_{{\mathbf{x}}_{n}}$ where $b_{\emptyset}\in\mathbf{K}_{\emptyset,\hat{t}_{\ast}}$ and $b_{{\mathbf{x}}_{i}}\in\mathbf{K}_{{\mathbf{x}}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$, such that
$\sum_{i:r\in{\mathbf{x}}_{i}}t_{\ast}(b_{{\mathbf{x}}_{i}})t_{r}(X_{r})=(0)\text{ for all }r\in\emptyset\cup\left(\bigcup_{i=1}^{n}{\mathbf{x}}_{i}\right)=\bigcup_{i=1}^{n}{\mathbf{x}}_{i}.$
Then the previous arguments show that $q_{\operatorname{c}}^{\rm{N}}(b_{{\mathbf{x}}_{i}})=0$ for all $i=1,\dots,n$. Moreover by Proposition 3.5 for $\tilde{t}$ we also have that $q_{\operatorname{c}}^{\rm{N}}(b_{\emptyset})=0$. Hence we get that
$t_{\ast}(b_{\emptyset})+\sum_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=\dot{t}(q_{\operatorname{c}}^{\rm{N}}(b_{\emptyset}))+\sum_{i=1}^{n}\dot{t}(q_{\operatorname{c}}^{\rm{N}}(b_{\mathbf{x}_{i}}))=0,$
as required, and the proof is complete. ∎
As a corollary we have an alternative proof of [11, Proposition 4.3].
Corollary 3.7.
Let $P$ be a unital right LCM semigroup that is a subsemigroup of a discrete group $G$ and let $X$ be a compactly aligned product system over $P$ in the sense of Fowler. Then ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)$ and ${\mathcal{N}}{\mathcal{T}}(X)$ are canonically isomorphic. In particular, a representation of $X$ is Fock covariant if and only if it is Nica covariant.
Proof..
This is immediate by Proposition 3.6 since the universal Fock covariant and the universal Nica covariant representations are both equivariant and injective. ∎
3.3. The product system of a semigroup
One class of product systems of particular interest arises from the semigroup representations of $P$ itself. Fock covariance in this case is described by Laca and Sehnem [34]. We consider $P$ to be a unital subsemigroup of a discrete group $G$. Following the notation of [34], we write $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ for the words of even length where $p_{1},q_{1},\dots,p_{n},q_{n}\in P$. A word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ is called neutral if $p_{1}^{-1}q_{1}\cdots p_{n}^{-1}q_{n}=e$. For a word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ we write
$K(\alpha):=q_{n}^{-1}p_{n}\dots q_{1}^{-1}p_{1}P$
for the induced constructible ideal in ${\mathcal{J}}$. For a map $w\colon P\to{\mathcal{B}}(H)$ we write
$\dot{w}_{\alpha}:=w_{p_{1}}^{*}w_{q_{1}}\cdots w_{p_{n}}^{*}w_{q_{n}}\in{\mathcal{B}}(H).$
It follows that
$\dot{w}_{\alpha}\dot{w}_{\beta}=\dot{w}_{\alpha\beta},\text{ for all words $\alpha,\beta$,}$
where $\alpha\beta$ denotes the concatenation of the words $\alpha$ and $\beta$.
A map $w\colon P\to{\mathcal{B}}(H)$ is called a representation of $P$ if it is a semigroup homomorphism. A representation $w$ of $P$ will be called equivariant if there exists a *-homomorphism $\delta$ of $\mathrm{C}^{*}(w)$ such that
$\delta\colon\mathrm{C}^{*}(w)\to\mathrm{C}^{*}(w)\otimes\mathrm{C}^{*}_{\max}(G);w_{p}\mapsto w_{p}\otimes u_{p}.$
It follows that $\delta$ is injective with a left inverse given by the map ${\operatorname{id}}\otimes\chi$. Moreover, it satisfies the coaction identity and hence $\mathrm{C}^{*}(w)$ admits a coaction by $G$. For simplicity, we will say that $w$ admits a coaction by $G$ if such a $\delta$ exists.
Remark 3.8.
Let $w\colon P\to{\mathcal{B}}(H)$ be an isometric representation of $P$ (i.e., every $w_{p}$ is an isometry) and let $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ where $K(\alpha)\not=\emptyset$. Since every $w_{p}$ is an isometry we have that
$w_{p}^{*}w_{pq}=w_{p}^{*}w_{p}w_{q}=w_{q},\text{ for all }p,q\in P.$
Therefore a variant of the proof of Proposition 2.3 yields
$\dot{w}_{\alpha}w_{r}=w_{p_{1}^{-1}q_{1}\cdots p_{n}^{-1}q_{n}r}\text{ for all}r\in K(\alpha).$
In particular, if $\alpha$ is neutral we obtain that
$\dot{w}_{\alpha}w_{r}=w_{r}\text{ for all }r\in K(\alpha).$
In [34, Proposition 3.2] it is shown that if a map $w\colon P\to{\mathcal{B}}(H)$ satisfies
  • (T1)
    $w_{e}=1$,
  • (T2)
    $\dot{w}_{\alpha}=0$ if $K(\alpha)=\emptyset$ for a neutral word $\alpha$, and
  • (T3)
    $\dot{w}_{\alpha}=\dot{w}_{\beta}$ if $K(\alpha)=K(\beta)$ for neutral words $\alpha$ and $\beta$,
then $w$ is an isometric representation of $P$, and the operators $\{\dot{w}_{\alpha}\}_{\alpha:\text{neutral}}$ are commuting projections. Let ${\mathcal{T}}(P)$ be the universal C*-algebra with respect to the isometric representations of $P$. By the universal property, the C*-algebra ${\mathcal{T}}(P)$ admits a coaction by $G$ and hence a topological C*-grading.
Let ${\mathcal{T}}_{u}(P)$ be the universal C*-algebra with respect to the maps $w\colon P\to{\mathcal{B}}(H)$ that satisfy conditions (T1)–(T3) and the additional condition:
  • (T4)
    $\prod_{\beta\in F}(\dot{w}_{\alpha}-\dot{w}_{\beta})=0$ whenever $\alpha$ is a neutral word, $F$ is a finite set of neutral words and $K(\alpha)=\bigcup_{\beta\in F}K(\beta)$.
Since conditions (T1)–(T4) are graded with respect to the grading of ${\mathcal{T}}(P)$, the C*-algebra ${\mathcal{T}}_{u}(P)$ is a quotient of ${\mathcal{T}}(P)$ by an induced ideal, and therefore ${\mathcal{T}}_{u}(P)$ admits a coaction by $G$. By using similar arguments as in Proposition 2.7 it can be proved that
$\mathrm{C}^{*}_{\max}\left(\Big{\{}[{\mathcal{T}}(P)]_{g}\Big{\}}_{g\in G}\right)\simeq{\mathcal{T}}(P),$
and hence combining with Proposition 2.2 we obtain that
$\mathrm{C}^{*}_{\max}\left(\Big{\{}[{\mathcal{T}}_{u}(P)]_{g}\Big{\}}_{g\in G}\right)\simeq{\mathcal{T}}_{u}(P).$
In [34, Corollary 3.19, Corollary 3.20] it is shown that there is a *-epimorphism ${\mathcal{T}}_{u}(P)\to\mathrm{C}^{*}_{\lambda}(P)$ that is equivariant and injective on the fixed point algebra. Hence ${\mathcal{T}}_{u}(P)$ coincides with the universal C*-algebra of the Fell bundle induced in $\mathrm{C}^{*}_{\lambda}(P)$; see also [22, Remark 3.15].
Fix $\{v_{p}\}_{p\in P}$ such that ${\mathcal{T}}(P)=\mathrm{C}^{*}(v)$. We can then define the concrete product system $X_{P}$ in ${\mathcal{T}}(P)$ by
$X_{P,p}:=\mathbb{C}v_{p}\text{ for all }p\in P.$
We say that a representation $t$ of $X_{P}$ is unital if it satisfies $t_{e}(v_{e})=1$. Note that a unital representation is automatically injective. If $t\colon X_{P}\to{\mathcal{B}}(H)$ is a non-zero representation, then $t_{e}(v_{e})$ is a projection that commutes with $\mathrm{C}^{*}(t)$ and hence $K:=t_{e}(v_{e})H$ is reducing for $\mathrm{C}^{*}(t)$. Therefore by compressing on $K$ we can consider $t$ to be a unital representation of $X_{P}$. Note that the Fock space of $X_{P}$ is unitarily equivalent to $\ell^{2}(P)$ by the unitary
(3.4) $W_{P}\colon{\mathcal{F}}X_{P}\to\ell^{2}(P);v_{p}\mapsto\delta_{p}.$
It follows that
(3.5) ${\mathcal{T}}_{\lambda}(X_{P})\simeq\mathrm{C}^{*}_{\lambda}(P)$
by the canonical *-isomorphism induced by $\operatorname{ad}_{W_{P}}$, since $W_{P}\lambda_{p}(v_{p})W_{P}^{*}=V_{p}$ for all $p\in P$. Since this *-isomorphism is canonical we have that the Fell bundles induced in ${\mathcal{T}}_{\lambda}(X_{P})$ and $\mathrm{C}^{*}_{\lambda}(P)$ are isomorphic, and therefore we also have that
(3.6) ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X_{P})\simeq{\mathcal{T}}_{u}(P).$
More generally, the association
$t=\{t_{p}\}_{p\in P}\mapsto w_{t}=\{t_{p}(v_{p})\}_{p\in P}$
induces a bijection between the unital (equivariant) representations of $X_{P}$ and the unital (resp. equivariant) isometric representations of $P$. Hence we have that
(3.7) ${\mathcal{T}}(X_{P})\simeq{\mathcal{T}}(P)$
by a canonical *-isomorphism. Note that Laca and Sehnem [34, Definition 3.6] coin ${\mathcal{T}}_{u}(P)$ as the Toeplitz algebra of $P$; however we will not use this terminology, as ${\mathcal{T}}_{u}(P)$ is not ${\mathcal{T}}(X_{P})$. Below we give an alternative proof of (3.6) that squares with our characterisation of Fock covariance.
Proposition 3.9.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X_{P}$ be the induced product system in ${\mathcal{T}}(P)$. Then the association
$t=\{t_{p}\}_{p\in P}\mapsto w_{t}=\{t_{p}(v_{p})\}_{p\in P}$
defines a bijection between the unital equivariant Fock covariant representations of $X_{P}$ and the equivariant representations of $P$ that satisfy conditions (T1)–(T4).
Proof..
To fix notation, let ${\mathcal{T}}(P)=\mathrm{C}^{*}(v)$, ${\mathcal{T}}(X_{P})=\mathrm{C}^{*}(\hat{t})$ and ${\mathcal{T}}_{u}(P)=\mathrm{C}^{*}(\tilde{u})$. It is readily verified that the association $t\mapsto w_{t}$ is a bijection between the unital equivariant representations of $X_{P}$ and the unital equivariant isometric representations of $P$. When it is clear from the context we will simply write $w$ instead of $w_{t}$. With this notation we have that
$t_{\ast}\left(\hat{t}_{p_{1}}(v_{p_{1}})^{*}\hat{t}_{q_{1}}(v_{q_{1}})\cdots\hat{t}_{p_{n}}(v_{p_{n}})^{*}\hat{t}_{q_{n}}(v_{q_{n}})\right)=\dot{w}_{\alpha}$
for every word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ where $p_{1},q_{1},\dots,p_{n},q_{n}\in P$.
First suppose that $t$ is a unital equivariant Fock covariant representation of $X_{P}$, and let $w$ be the associated unital equivariant representation of $P$. Then $t$ satisfies conditions (i) and (ii) of Theorem 3.2. For condition (T1), we have that $t_{e}(v_{e})$ is the unit, and thus $w_{e}=t_{e}(v_{e})=1$.
For condition (T2), let $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ be a neutral word such that $K(\alpha)=\emptyset$. We then have that
$\dot{w}_{\alpha}=t_{\ast}\left(\hat{t}_{p_{1}}(v_{p_{1}})^{*}\hat{t}_{q_{1}}(v_{q_{1}})\cdots\hat{t}_{p_{n}}(v_{p_{n}})^{*}\hat{t}_{q_{n}}(v_{q_{n}})\right)\in\mathbf{K}_{\emptyset,t_{\ast}}=(0),$
as required.
For condition (T3), let $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ and $\beta=(r_{1},s_{1},\dots,r_{m},s_{m})$ be neutral words such that $K(\alpha)=K(\beta)$. For brevity, set ${\mathbf{x}}:=K(\alpha)=K(\beta)$ and
$b_{\alpha}:=\hat{t}_{p_{1}}(v_{p_{1}})^{*}\hat{t}_{q_{1}}(v_{q_{1}})\cdots\hat{t}_{p_{n}}(v_{p_{n}})^{*}\hat{t}_{q_{n}}(v_{q_{n}})\quad\text{and}\quad b_{\beta}:=\hat{t}_{r_{1}}(v_{r_{1}})^{*}\hat{t}_{s_{1}}(v_{s_{1}})\cdots\hat{t}_{r_{m}}(v_{r_{m}})^{*}\hat{t}_{s_{m}}(v_{s_{m}}).$
Let $r\in\mathbf{x}$ and $\xi_{r}=\mu v_{r}\in X_{P,r}$ with $\mu\in\mathbb{C}$. Then Proposition 2.5, along with that $\dot{V}_{\alpha}=\dot{V}_{\beta}$, yields
$t_{\ast}(b_{\alpha})t_{r}(\xi_{r})$ $=t_{r}(\lambda_{\ast}(b_{\alpha})\xi_{r})=\mu t_{r}(W_{P}^{*}\dot{V}_{\alpha}\delta_{r})=\mu t_{r}(W_{P}^{*}\dot{V}_{\beta}\delta_{r})=t_{r}(\lambda_{\ast}(b_{\beta})\xi_{r})=t_{\ast}(b_{\beta})t_{r}(\xi_{r}),$
for the unitary $W_{P}$ of (3.4). Hence, we have that
$(t_{\ast}(b_{\alpha})-t_{\ast}(b_{\beta}))t_{r}(\xi_{r})=0\text{ for all }\xi_{r}\in X_{P,r},r\in{\mathbf{x}}.$
Thus applying condition (ii) for $t$ implies that $t_{\ast}(b_{\alpha})-t_{\ast}(b_{\beta})=0$, and therefore
$\dot{w}_{\alpha}=t_{\ast}(b_{\alpha})=t_{\ast}(b_{\beta})=\dot{w}_{\beta},$
as required.
For condition (T4), let $F$ be a finite set of neutral words, and let $\alpha$ be a neutral word such that $K(\alpha)=\bigcup_{\beta\in F}K(\beta)$. Let ${\mathcal{F}}$ be the $\cap$-closure of $\{K(\beta)\mid\beta\in F\}\cup\{K(\alpha)\}\cup\{\emptyset\}$. For each $D\subseteq F$ we write $\beta_{D}$ for the neutral word that arises by concatenating the words $\beta\in D$ in some order. As the induced constructible ideal does not depend on the order of the concatenation we choose (being on neutral words), we have that
$K(\beta_{D})=\bigcap_{\beta\in D}K(\beta),$
and that
=${\mathcal{F}}=\{K(\beta_{D}):\emptyset\not=D\subseteq F\}\cup\{K(\alpha)\}\cup\{\emptyset\}.$
In particular we have that
$\dot{w}_{\beta_{D}}=\prod_{\beta\in D}\dot{w}_{\beta}.$
For each $D\subseteq F$, let $b_{D}\in\mathbf{K}_{K(\beta_{D}),\hat{t}_{\ast}}$ and $b_{\alpha}\in\mathbf{K}_{K(\alpha),\hat{t}_{\ast}}$ such that
$t_{\ast}(b_{D})=\dot{w}_{\beta_{D}}\quad\text{and}\quad t_{\ast}(b_{\alpha})=\dot{w}_{\alpha}.$
Then, for $r\in K(\alpha)$ and $\mu v_{r}\in X_{P,r}$, we have that
$\left(t_{\ast}(b_{\alpha})+\sum_{\begin{subarray}{c}\emptyset\neq D\subseteq F:\\ r\in K(\beta_{D})\end{subarray}}(-1)^{|D|}t_{\ast}(b_{\beta_{D}})\right)t_{r}(\mu v_{r})$ $=\left(\dot{w}_{\alpha}+\sum_{\begin{subarray}{c}\emptyset\neq D\subseteq F:\\ r\in K(\beta_{D})\end{subarray}}(-1)^{|D|}\dot{w}_{\beta_{D}}\right)\mu w_{r}$
$=\left(1+\sum_{\begin{subarray}{c}\emptyset\neq D\subseteq F:\\ r\in\bigcap\limits_{\beta\in D}K(\beta)\end{subarray}}(-1)^{|D|}\right)\mu w_{r}$
$=\left(1+\sum_{\emptyset\neq D\subseteq\{\beta\in F\mid r\in K(\beta)\}}(-1)^{|D|}\right)\mu w_{r}=0,$
where we used that
$\dot{w}_{\alpha}w_{r}=w_{r}\quad\text{and}\quad\dot{w}_{\beta_{D}}w_{r}=w_{r}\text{ when }r\in K(\beta_{D}).$
Note also that in the third equality we used that
$\{\emptyset\neq D\subseteq F\mid r\in\bigcap\limits_{\beta\in D}K(\beta)\}={\mathcal{P}}(\{\beta\in F\mid r\in K(\beta)\})\setminus\{\emptyset\},$
and that these are non-empty sets since $r\in K(\alpha)=\bigcup_{\beta\in F}K(\beta)$. Condition (ii) for $t$ then implies that
$t_{\ast}(b_{\alpha})+\sum_{\emptyset\neq D\subseteq F}(-1)^{|D|}t_{\ast}(b_{\beta_{D}})=0,$
and therefore we have that
$\prod_{\beta\in F}(\dot{w}_{\alpha}-\dot{w}_{\beta})$ $=\dot{w}_{\alpha}+\sum_{\emptyset\neq D\subseteq F}(-1)^{|D|}\prod_{\beta\in D}\dot{w}_{\beta}$
$=\dot{w}_{\alpha}+\sum_{\emptyset\neq D\subseteq F}(-1)^{|D|}\dot{w}_{\beta_{D}}=t_{\ast}(b_{\alpha})+\sum_{\emptyset\neq D\subseteq F}(-1)^{|D|}t_{\ast}(b_{\beta_{D}})=0,$
as required.
For the reverse implication, suppose that $w$ is an equivariant representation of $P$ that satisfies conditions (T1)–(T4) and let $t$ be the unital equivariant representation of $X_{P}$ associated with $w$. We will show that $t$ satisfies conditions (i) and (ii) of Theorem 3.2.
Let $\mathbf{x}$ be in ${\mathcal{J}}$ and pick a neutral word $\alpha=(p_{1},q_{1},\dots,p_{n},q_{n})$ such that $\mathbf{x}=K(\alpha)$. Then condition (T3) implies that
$t_{\ast}(\mathbf{K}_{\mathbf{x},\hat{t}_{\ast}})=\text{span}\{\dot{w}_{\alpha}\}.$
In particular, combining (T2) with (T3) yields that $t_{\ast}(\mathbf{K}_{\emptyset,\hat{t}_{\ast}})=(0)$, and thus $t$ satisfies condition (i) of Theorem 3.2.
For condition (ii), let =${\mathcal{F}}=\{{\mathbf{x}_{1},\dots,\mathbf{x}_{n}}\}$ be a finite $\cap$-closed subset of ${\mathcal{J}}$ such that $\bigcup_{i=1}^{n}\mathbf{x}_{i}\neq\emptyset$, and let $b_{\mathbf{x}_{i}}\in\mathbf{K}_{\mathbf{x}_{i},\hat{t}_{\ast}}$ for $i=1,\dots,n$. Suppose that
$\sum_{i:r\in\mathbf{x}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{r}(X_{P,r})=0\text{ for all }r\in\bigcup_{i=1}^{n}\mathbf{x}_{i},$
and we will show that $\sum_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}})=0$. For notational convenience, set
$b:=\sum_{i=1}^{n}t_{\ast}(b_{\mathbf{x}_{i}}).$
First notice that $t_{\ast}(\mathbf{K}_{\emptyset,\hat{t}_{\ast}})=(0)$ as $t$ satisfies condition (i), therefore without loss of generality we may assume that ${\mathbf{x}}_{i}\neq\emptyset$ for all $i=1,\dots,n$. Next let
$F:=\{\alpha_{1},\dots,\alpha_{n}\}$
be a set of neutral words such that $K(\alpha_{i})=\mathbf{x}_{i}$ for $i=1,\dots,n$. Moreover, from the comments above we may pick $\mu_{i}\in\mathbb{C}$ such that $t_{\ast}(b_{\mathbf{x}_{i}})=\mu_{i}\dot{w}_{\alpha_{i}}$ for $i=1,\dots,n$. As the projections $\dot{w}_{\alpha_{i}}$ commute, by the Gel’fand-Naimark Theorem, and considering a unit decomposition (see also [34, Lemma 3.13]), there exists a subset $B$ of $F$ such that
$\|b\|=\|Q_{B}b\|\quad\text{for}\quad Q_{B}:=\prod_{i:\alpha_{i}\in B}\dot{w}_{\alpha_{i}}\prod_{j:\alpha_{j}\not\in B}(1-\dot{w}_{\alpha_{j}}).$
Since $b=\sum_{i=1}^{n}\mu_{i}\dot{w}_{\alpha_{i}}$, by the form of the projection $Q_{B}$ we have that
$\|b\|=\|Q_{B}b\|=\|\sum_{i:\alpha_{i}\in B}\mu_{i}Q_{B}\|=|\sum_{i:\alpha_{i}\in B}\mu_{i}|.$
We consider the following cases.
Case 1. Suppose that
$\bigcap_{i:\alpha_{i}\in B}K(\alpha_{i})\subseteq\bigcup_{j:\alpha_{j}\not\in B}K(\alpha_{j}).$
Then we have that
$\bigcap_{i:\alpha_{i}\in B}K(\alpha_{i})=\bigcup_{j:\alpha_{j}\not\in B}\left(K(\alpha_{j})\cap\left(\bigcap_{i:\alpha_{i}\in B}K(\alpha_{i})\right)\right).$
Applying (T4) for
$K(\alpha)=\bigcap_{i:\alpha_{i}\in B}K(\alpha_{i}),$
where $\alpha$ is a concatenation of the words $\alpha_{i}\in B$, and
$K(\beta_{j})=K(\alpha_{j})\cap\left(\bigcap_{i:\alpha_{i}\in B}K(\alpha_{i})\right)\text{ such that }\alpha_{j}\notin B,$
where $\beta_{j}$ is a concatenation of the words $\alpha_{i}\in B$ and $\alpha_{j}$, gives that
$\prod_{j:\alpha_{j}\notin B}(\dot{w}_{\alpha}-\dot{w}_{\beta_{j}})=0.$
By construction, for any $j$ with $\alpha_{j}\notin B$ we have that
$\dot{w}_{\beta_{j}}\leq\dot{w}_{\alpha_{j}}\quad\text{and}\quad\dot{w}_{\beta_{j}}\leq\dot{w}_{\alpha}=\prod_{i:\alpha_{i}\in B}\dot{w}_{\alpha_{i}}.$
Therefore we have that
$Q_{B}=\dot{w}_{\alpha}\prod_{j:\alpha_{j}\not\in B}(1-\dot{w}_{\alpha_{j}})\leq\dot{w}_{\alpha}\prod_{j:\alpha_{j}\notin B}(1-\dot{w}_{\beta_{j}})=\prod_{j:\alpha_{j}\notin B}(\dot{w}_{\alpha}-\dot{w}_{\beta_{j}})=0,$
and thus $b=0$.
Case 2. Suppose that
$\bigcap_{i:\alpha_{i}\in B}K(\alpha_{i})\not\subseteq\bigcup_{j:\alpha_{j}\not\in B}K(\alpha_{j}).$
Then there exists a $p\in\bigcap_{i:\alpha_{i}\in B}K(\alpha_{i})\subseteq\bigcup_{i=1}^{n}K(\alpha_{i})$ such that $p\not\in\bigcup_{j:\alpha_{j}\not\in B}K(\alpha_{j})$. A moment’s thought shows that this dichotomy implies that
$\{i\mid\alpha_{i}\in B\}=\{i\mid p\in K(\alpha_{i})\}.$
Since $w_{p}$ is an isometry and $\dot{w}_{\alpha_{i}}w_{p}=w_{p}$ for all $i$ with $p\in K(\alpha_{i})$, we conclude that
$\|b\|$ $=|\sum_{i:\alpha_{i}\in B}\mu_{i}|=|\sum_{i:p\in K(\alpha_{i})}\mu_{i}|=\|\sum_{i:p\in K(\alpha_{i})}\mu_{i}\dot{w}_{\alpha_{i}}w_{p}\|=\|\sum_{i:p\in{\mathbf{x}}_{i}}t_{\ast}(b_{\mathbf{x}_{i}})t_{p}(v_{p})\|=0,$
where we applied the assumption for $p\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$ in the last equality. Thus $b=0$ in this case as well, and the proof is complete. ∎
Corollary 3.10.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X_{P}$ be the associated product system. Then ${\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X_{P})$ and ${\mathcal{T}}_{u}(P)$ are canonically isomorphic. In particular, a representation of $X_{P}$ is Fock covariant if and only if the associated semigroup representation of $P$ satisfies conditions (T1)–(T4).
Proof..
This is immediate by Proposition 3.9 since the universal Fock covariant representation of $X_{P}$ is equivariant and the left action on $X_{P}$ is unital. ∎
4. The reduced Hao–Ng isomorphism problem
In this section we provide the affirmative solution to the reduced Hao–Ng isomorphism problem for generalised gauge actions by discrete groups. In order to put the problem into context we require some elements from strong covariant representations and tensor algebras. We will also require some elements from the theory of nonselfadjoint operator algebras.
4.1. Operator algebras
By an operator algebra ${\mathfrak{A}}$ we will mean a norm-closed subalgebra of some ${\mathcal{B}}(H)$. Every operator algebra attains a C*-cover, i.e., a completely isometric homomorphism $\iota\colon{\mathfrak{A}}\to{\mathcal{C}}$ such that ${\mathcal{C}}=\mathrm{C}^{*}(\iota({\mathfrak{A}}))$. The C*-envelope $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ is the co-universal C*-cover of ${\mathfrak{A}}$, i.e., there exists a completely isometric homomorphism $\iota\colon{\mathfrak{A}}\to\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ such that for any other C*-cover $j\colon{\mathfrak{A}}\to{\mathcal{C}}$ there exists a unique *-epimorphism $\Phi\colon{\mathcal{C}}\to\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ such that $\Phi\circ j=\iota$. The reader is addressed to [5, 44] for the full details.
The existence of the C*-envelope was established by Hamana [19] through the existence of the injective envelope. An independent proof was established by Dritschel and McCullough [14] through the existence of maximal dilations. Recall that a homomorphism $\phi\colon{\mathfrak{A}}\to{\mathcal{B}}(K)$ is called a dilation of $\iota\colon{\mathfrak{A}}\to{\mathcal{B}}(H)$ if $H\subseteq K$ and $\iota(\cdot)=P_{H}\phi(\cdot)|_{H}$. A dilation is called maximal if it only attains trivial dilations, i.e., dilations by orthogonal summands. As an application of the existence of maximal dilations, Dritschel and McCullough [14] show that the C*-envelope of ${\mathfrak{A}}$ is the C*-algebra generated by a maximal dilation. Arveson [3] has shown that a representation of ${\mathfrak{A}}$ is maximal if and only if it has a unique extension to a completely positive map on $\mathrm{C}^{*}({\mathfrak{A}})$ that is a *-homomorphism. The proof in [3] refers to operator systems, but it can be adapted to the operator algebras category; the reader is directed to [13] for the full details.
4.2. Strong covariant representations
We review the key elements of the strong covariant C*-algebra of a product system established by Sehnem [47, 48], and its Fell bundle considered by Dor-On, the first author, Katsoulis, Laca and Li [11].
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Set $A:=X_{e}$, and for each $p\in P$ let $\varphi_{p}\colon A\to{\mathcal{L}}(X_{p})$ be the *-homomorphism that implements the left action of $A$ on the C*-correspondence $X_{p}$. For a finite set $F\subseteq G$ let
$K_{F}:=\bigcap_{g\in F}gP.$
For $r\in P$ and $g\in F$ define the ideal of $A$ given by
$I_{r^{-1}K_{\{r,g\}}}:=\begin{cases}\bigcap\limits_{s\in K_{\{r,g\}}}\ker\varphi_{r^{-1}s}&\text{if }K_{\{r,g\}}\neq\emptyset\text{ and }r\notin K_{\{r,g\}},\\ \hskip 14.22636ptA&\text{otherwise},\end{cases}$
and set
$I_{r^{-1}(r\vee F)}:=\bigcap_{g\in F}I_{r^{-1}K_{\{r,g\}}}.$
We have that $I_{r^{-1}(r\vee F)}=I_{(pr)^{-1}(pr\vee pF)}$ for all $r,p\in P$, and $I_{r^{-1}(r\vee F)}=I_{(s^{-1}r)^{-1}(s^{-1}r\vee s^{-1}F)}$ for all $r\in sP$. Moreover
$I_{r^{-1}(r\vee F_{1})}\subseteq I_{r^{-1}(r\vee F_{2})}\;\text{ when }\;F_{1}\supseteq F_{2}.$
We declare $K_{\emptyset}=\emptyset$ and $I_{r^{-1}(r\vee\emptyset)}=A$. For a finite set $F\subseteq G$, let the C*-correspondences
$X_{F}:=\operatornamewithlimits{\sum^{\oplus}}_{r\in P}X_{r}I_{r^{-1}(r\vee F)}\quad\text{and}\quad X_{F}^{+}:=\operatornamewithlimits{\sum^{\oplus}}_{g\in G}X_{gF}.$
We declare $X_{\emptyset}=X_{\emptyset}^{+}={\mathcal{F}}X$. Each $X_{F}^{+}$ is reducing for the coaction
$\delta\colon{\mathcal{T}}_{\lambda}(X)\longrightarrow{\mathcal{T}}_{\lambda}(X)\otimes\mathrm{C}^{*}_{\max}(G);\lambda_{p}(\xi_{p})\mapsto\lambda_{p}(\xi_{p})\otimes u_{p},$
giving rise to a *-representation
$\Phi_{F}\colon{\mathcal{T}}_{\lambda}(X)\longrightarrow{\mathcal{L}}(X_{F}^{+});\lambda_{p}(\xi_{p})\mapsto(\lambda_{p}(\xi_{p})\otimes u_{p})|_{X_{F}^{+}}.$
Here we make the identification
$X_{gF}\longrightarrow X_{gF}\otimes\delta_{g};\xi_{r}a_{r}\mapsto\xi_{r}a_{r}\otimes\delta_{g},\text{ for }\xi_{r}\in X_{r},a_{r}\in I_{r^{-1}(r\vee gF)}.$
Moreover $X_{F}\subseteq X_{F}^{+}$ is reducing for $[{\mathcal{T}}(X)]_{e}$ and so we obtain the representation
$\bigoplus\limits_{\textup{fin }F\subseteq G}\Phi_{F}(\cdot)|_{X_{F}}\colon[{\mathcal{T}}_{\lambda}(X)]_{e}\longrightarrow\prod\limits_{\text{fin }F\subseteq G}{\mathcal{L}}(X_{F}).$
We fix the ideal ${\mathcal{I}}_{\operatorname{sc},e}$ in $[{\mathcal{T}}(X)]_{e}$ by using the corona universe, namely
$b\in{\mathcal{I}}_{\operatorname{sc},e}\quad\text{if and only if}\quad\bigoplus\limits_{\textup{fin }F\subseteq G}\Phi_{F}(\lambda_{\ast}(b))|_{X_{F}}\in c_{0}\left({\mathcal{L}}(X_{F})\mid\textup{fin }F\subseteq G\right).$
We write $A\times_{X}P$ for the equivariant quotient of ${\mathcal{T}}(X)$ by the induced ideal ${\mathcal{I}}_{\operatorname{sc}}:=\left\langle{\mathcal{I}}_{\operatorname{sc},e}\right\rangle$. It is shown in [47] that this construction does not depend on $G$. Moreover, we define the strong covariant bundle of $X$ to be the Fell bundle
${\mathcal{S}}{\mathcal{C}}X:=\Big{\{}[A\times_{X}P]_{g}\Big{\}}_{g\in G},$
given by the coaction by $G$ on $A\times_{X}P$. A representation of $X$ that promotes to a representation of ${\mathcal{S}}{\mathcal{C}}X$ will be called a strong covariant representation of $X$. By combining Proposition 2.2 with Proposition 2.7 we have that $A\times_{X}P$ is the universal C*-algebra of ${\mathcal{S}}{\mathcal{C}}X$. We further write $A\times_{X,\lambda}P$ for the reduced C*-algebra of ${\mathcal{S}}{\mathcal{C}}X$. As the strong covariant relations are graded and induced by representations on ${\mathcal{T}}_{\lambda}(X)$ it follows that a strong covariant representation is Fock covariant.
For a product system $X$ we will write ${\mathcal{T}}_{\lambda}(X)^{+}$ for the tensor algebra of $X$, i.e., for the norm-closed subalgebra generated by $\{\lambda_{p}(X_{p})\}_{p\in P}$ inside ${\mathcal{T}}_{\lambda}(X)$. We will require the following consequence of [48, Corollary 3.5]. We note that although only product systems in the sense of Fowler are considered in [48], the arguments apply in our setting as well.
Corollary 4.1.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. If $t$ is a Fock covariant injective representation of $X$ and $t$ admits a coaction by $G$ which is normal, then the map
${\mathcal{T}}_{\lambda}(X)^{+}\to\overline{\operatorname{alg}}\{t_{p}(X_{p})\mid p\in P\};\lambda_{p}(\xi_{p})\mapsto t_{p}(\xi_{p}),$
is a (well-defined) completely isometric isomorphism.
Proof..
Let the canonical *-epimorphism $\Phi\colon{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)\to\mathrm{C}^{*}(t)$. Since $t$ is equivariant, then $\Phi$ induces a *-epimorphism between the reduced C*-algebras of the induced Fell bundles. However, since the coaction on $\mathrm{C}^{*}(t)$ is normal, it follows that $\mathrm{C}^{*}(t)$ coincides with the reduced C*-algebra of its induced Fell bundle. Hence there exists a canonical *-epimorphism $\Phi_{\lambda}\colon{\mathcal{T}}_{\lambda}(X)\to\mathrm{C}^{*}(t)$, which is injective on $A$ since $t$ is injective. By [48, Corollary 3.5] it follows that the restriction of $\Phi_{\lambda}$ on ${\mathcal{T}}_{\lambda}(X)^{+}$ is completely isometric. ∎
Since $A\times_{X,\lambda}P$ satisfies the conditions of Corollary 4.1, it is a C*-cover of ${\mathcal{T}}_{\lambda}(X)^{+}$. In [48, Theorem 5.1], Sehnem establishes that $A\times_{X,\lambda}P$ is the C*-envelope of ${\mathcal{T}}_{\lambda}(X)^{+}$.
4.3. Crossed products
The theory of crossed products of C*-algebras is well-known, see for example [50]. Katsoulis and Ramsey [27] have extended this to group actions over possibly nonselfadjoint operator algebras. Here we will comment only on the parts that are relevant to the reduced Hao–Ng isomorphism problem.
Let ${\mathfrak{H}}$ be a locally compact group that acts on an operator algebra ${\mathfrak{A}}$ by completely isometric automorphisms. The action of ${\mathfrak{H}}$ then extends also to a group action $\dot{\alpha}$ of the C*-envelope of ${\mathfrak{A}}$, and we can form the reduced crossed product $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$. By considering the copy of ${\mathfrak{A}}$ inside $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$, the reduced crossed product ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is defined as the norm-closed subalgebra of $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$ generated by the ${\mathfrak{A}}$-valued functions [27, Definition 3.17]. One of the main questions in the theory is whether this inclusion induces a canonical *-isomorphism with the C*-envelope of ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$, i.e., whether it holds that
(4.1) $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}})\stackrel{{\scriptstyle?}}{{\simeq}}\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}.$
When ${\mathfrak{A}}$ admits a contractive approximate identity, this has been answered to the affirmative when ${\mathfrak{H}}$ is discrete [24], and when ${\mathfrak{H}}$ is abelian [27]. By using the maximal representations of [14] we can remove the contractive approximate identity hypothesis when ${\mathfrak{H}}$ is discrete.
Proposition 4.2.
Let ${\mathfrak{H}}$ be a discrete group acting by $\alpha$ on an operator algebra ${\mathfrak{A}}$. Then
$\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}})\simeq\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}.$
Proof..
We identify ${\mathfrak{A}}$ with its copy in $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$. Let $\pi\colon{\mathfrak{A}}\to{\mathcal{B}}(H)$ be a maximal completely isometric representation of ${\mathfrak{A}}$. Then $\pi$ has a unique extension to a faithful *-representation on $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$, denoted by the same symbol. Therefore $\overline{\pi}\rtimes U$ is a faithful *-representation of $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$, and thus its restriction on ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is completely isometric. It suffices to show that $\overline{\pi}\rtimes U$ is maximal on ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$, as then it will follow that
$\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}})\simeq\mathrm{C}^{*}(\overline{\pi}\rtimes U)\simeq\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$
by *-isomorphisms fixing ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$.
Towards this end, let $\rho\colon{\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}\to{\mathcal{B}}(K)$ be a maximal dilation of $\overline{\pi}\rtimes U|_{{\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}}$, and let us denote by the same symbol the unique extension of $\rho$ to a *-representation on
$\mathrm{C}^{*}({\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}})=\overline{\operatorname{span}}\{\overline{\pi}(c)U_{{\mathfrak{h}}}\mid c\in\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}}),{\mathfrak{h}}\in{\mathfrak{H}}\}.$
By Arveson’s Extension Principle, each $\pi\circ\alpha_{\mathfrak{h}}$ is a maximal representation of ${\mathfrak{A}}$, and it is a standard argument that $\overline{\pi}$ is maximal as the discrete direct sum of maximal representations, e.g., see [4, Proposition 4.4]. Therefore $\rho$ on $\overline{\pi}({\mathfrak{A}})$ takes up the form
$\rho(\overline{\pi}(a))=\begin{bmatrix}\overline{\pi}(a)&0\\ 0&\sigma(a)\end{bmatrix}\text{ for all }a\in{\mathfrak{A}},$
for a representation $\sigma$ of ${\mathfrak{A}}$. Next consider $\overline{\pi}(a)U_{{\mathfrak{h}}}$ for $a\in{\mathfrak{A}}$ and ${\mathfrak{h}}\in{\mathfrak{H}}$, and write
$\rho(\overline{\pi}(a)U_{\mathfrak{h}})=\begin{bmatrix}\overline{\pi}(a)U_{{\mathfrak{h}}}&x\\ y&z\end{bmatrix}.$
By using the unique extension property of $\rho$ we have that
$\begin{bmatrix}\overline{\pi}(a)\overline{\pi}(a)^{*}+xx^{*}&\ast\\ \ast&\ast\end{bmatrix}$ $=\rho(\overline{\pi}(a)U_{\mathfrak{h}})\rho(\overline{\pi}(a)U_{\mathfrak{h}})^{*}=\rho(\overline{\pi}(a)U_{\mathfrak{h}}U_{{\mathfrak{h}}}^{*}\overline{\pi}(a)^{*})$
$=\rho(\overline{\pi}(a)\overline{\pi}(a)^{*})=\rho(\overline{\pi}(a))\rho(\overline{\pi}(a))^{*}=\begin{bmatrix}\overline{\pi}(a)\overline{\pi}(a)^{*}&0\\ 0&\sigma(a)\sigma(a)^{*}\end{bmatrix}.$
By equating the $(1,1)$-entries we get that $x=0$. On the other hand, by using the covariance in the C*-crossed product we can write
$\overline{\pi}(\alpha_{\mathfrak{h}}^{-1}(a))^{*}\overline{\pi}(\alpha_{{\mathfrak{h}}}^{-1}(a))=\overline{\pi}(\alpha_{\mathfrak{h}}^{-1}(a))^{*}U_{{\mathfrak{h}}}^{*}U_{{\mathfrak{h}}}\overline{\pi}(\alpha_{{\mathfrak{h}}}^{-1}(a))=U_{\mathfrak{h}}^{*}\overline{\pi}(a)^{*}\overline{\pi}(a)U_{{\mathfrak{h}}}.$
Then a similar computation gives that
$\begin{bmatrix}\overline{\pi}(\alpha_{\mathfrak{h}}^{-1}(a))^{*}\overline{\pi}(\alpha_{{\mathfrak{h}}}^{-1}(a))+y^{*}y&\ast\\ \ast&\ast\end{bmatrix}$ $=\rho(\overline{\pi}(a)U_{\mathfrak{h}})^{*}\rho(\overline{\pi}(a)U_{\mathfrak{h}})=\rho\left(U_{\mathfrak{h}}^{*}\overline{\pi}(a)^{*}\overline{\pi}(a)U_{{\mathfrak{h}}}\right)$
$=\rho\left(\overline{\pi}(\alpha_{\mathfrak{h}}^{-1}(a))^{*}\overline{\pi}(\alpha_{{\mathfrak{h}}}^{-1}(a))\right)=\rho\left(\overline{\pi}(\alpha_{\mathfrak{h}}^{-1}(a))\right)^{*}\rho\left(\overline{\pi}(\alpha_{{\mathfrak{h}}}^{-1}(a))\right)$
$=\begin{bmatrix}\overline{\pi}(\alpha_{\mathfrak{h}}^{-1}(a))^{*}\overline{\pi}(\alpha_{\mathfrak{h}}^{-1}(a))&0\\ 0&\sigma(\alpha_{\mathfrak{h}}^{-1}(a))^{*}\sigma(\alpha_{\mathfrak{h}}^{-1}(a))\end{bmatrix}.$
By equating the $(1,1)$-entries we get that $y=0$. Therefore $\rho$ is a trivial dilation, as required. ∎
4.4. Hao–Ng
Let us return to the product system discussion. Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let ${\mathfrak{H}}$ be a discrete group, and let $\alpha$ be a generalised gauge action of ${\mathfrak{H}}$ on the Fock C*-algebra ${\mathcal{T}}_{\lambda}(X)$, i.e., every $\lambda_{p}(X_{p})$ is $\alpha$-invariant in the sense that
$\alpha_{{\mathfrak{h}}}(\lambda_{p}(X_{p}))=\lambda_{p}(X_{p})\text{ for all }p\in P\text{ and }{\mathfrak{h}}\in{\mathfrak{H}}.$
Let $\pi$ be a faithful *-representation of ${\mathcal{T}}_{\lambda}(X)$ in some ${\mathcal{B}}(H)$ and consider the faithful *-representation $\overline{\pi}\rtimes U$ on $\ell^{2}({\mathfrak{H}},H)$ that gives rise to ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Since $\alpha$ restricts to an action on ${\mathcal{T}}_{\lambda}(X)^{+}$, from [27, Corollary 3.16] we have a canonical completely isometric copy of the crossed product ${\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ in ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. We can now consider the family $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ of the subspaces
$(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})_{p}:=\overline{\operatorname{span}}\{\overline{\pi}(\lambda_{p}(\xi_{p}))U_{{\mathfrak{h}}}\mid\xi_{p}\in X_{p},{\mathfrak{h}}\in{\mathfrak{H}}\}\text{ for }p\in P.$
It is accustomed to use the short form $X_{p}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ instead of $(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})_{p}$. We will show that this construction gives rise to a product system.
Proposition 4.3.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\alpha$ be a generalised gauge action of a discrete group ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$. Then $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is a product system over $P$.
Proof..
For condition (i) we have by definition that
$X_{e}\rtimes_{\alpha,\lambda}{\mathfrak{H}}:=\overline{\operatorname{span}}\{\overline{\pi}(\lambda_{e}(\xi_{e}))U_{{\mathfrak{h}}}\mid\xi_{e}\in X_{e},{\mathfrak{h}}\in{\mathfrak{H}}\}\simeq A\rtimes_{\alpha,\lambda}{\mathfrak{H}},$
which is a C*-algebra. The other two conditions follow by the fact that $\alpha_{{\mathfrak{h}}}(\lambda_{p}(X_{p}))=\lambda_{p}(X_{p})$ for all $p\in P$ and ${\mathfrak{h}}\in{\mathfrak{H}}$, the covariant relations of the crossed product and that $\lambda$ is a product system representation. Indeed, for condition (ii) we have that
$\overline{\pi}(\lambda_{p}(X_{p}))U_{{\mathfrak{h}}_{1}}\overline{\pi}(\lambda_{q}(X_{q}))U_{{\mathfrak{h}}_{2}}$ $=\overline{\pi}\left(\lambda_{p}(X_{p})\alpha_{{\mathfrak{h}}_{1}}(\lambda_{q}(X_{q}))\right)U_{{\mathfrak{h}}_{1}{\mathfrak{h}}_{2}}$
$=\overline{\pi}\left(\lambda_{p}(X_{p})\lambda_{q}(X_{q})\right)U_{{\mathfrak{h}}_{1}{\mathfrak{h}}_{2}}\subseteq\overline{\pi}\left(\lambda_{pq}(X_{p}X_{q})\right)U_{{\mathfrak{h}}_{1}{\mathfrak{h}}_{2}}\subseteq X_{pq}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
By considering finite linear combinations and their norm-limits we obtain
$(X_{p}\rtimes_{\alpha,\lambda}{\mathfrak{H}})\cdot(X_{q}\rtimes_{\alpha,\lambda}{\mathfrak{H}})\subseteq X_{pq}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
For condition (iii) we have that
$\left(\overline{\pi}(\lambda_{p}(X_{p}))U_{{\mathfrak{h}}_{1}}\right)^{*}\overline{\pi}(\lambda_{pq}(X_{pq}))U_{{\mathfrak{h}}_{2}}$ $=\overline{\pi}\circ\alpha_{{\mathfrak{h}}_{1}}^{-1}\left(\lambda_{p}(X_{p})^{*}\lambda_{pq}(X_{pq})\right)U_{{\mathfrak{h}}_{1}^{-1}{\mathfrak{h}}_{2}}$
$\subseteq\overline{\pi}\left(\lambda_{q}(X_{q})\right)U_{{\mathfrak{h}}_{1}^{-1}{\mathfrak{h}}_{2}}\subseteq X_{q}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
By considering finite linear combinations and their norm-limits we obtain that
$(X_{p}\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{*}\cdot(X_{pq}\rtimes_{\alpha,\lambda}{\mathfrak{H}})\subseteq X_{q}\rtimes_{\alpha,\lambda}{\mathfrak{H}},$
and the proof is complete. ∎
In order to facilitate comparisons we will use the superscript $\rtimes$ for the representations of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. That is, we write $\lambda^{\rtimes}$ for the Fock representation of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$, and we fix a representation $\hat{t}^{\rtimes}$ of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ such that
${\mathcal{T}}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})=\mathrm{C}^{*}(\hat{t}^{\rtimes}).$
We will also write $\iota^{\rtimes}$ for the identity representation $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\hookrightarrow{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$.
Due to the properties of the generalised gauge action we have a canonical identification of the $\mathbf{K}_{\bullet}$-cores. Note that, since every $\lambda_{p}(X_{p})$ is $\alpha$-invariant, then so is every C*-subalgebra $\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}$ for ${\mathbf{x}}\in{\mathcal{J}}$. By construction we have that
$\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}\subseteq{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}\text{ for all }{\mathbf{x}}\in{\mathcal{J}}.$
Proposition 4.4.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\alpha$ be a generalised gauge action of a discrete group ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$ and let the identity representation $\iota^{\rtimes}\colon X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\to{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. If ${\mathbf{x}}\in{\mathcal{J}}$, then
$\iota^{\rtimes}_{\ast}(\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}})=\mathbf{K}_{{\mathbf{x}},\iota^{\rtimes}_{\ast}}=\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
Consequently, if $E_{\mathfrak{H}}\colon{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}\to{\mathcal{T}}_{\lambda}(X)$ is the faithful conditional expectation of the reduced crossed product, then
$E_{{\mathfrak{H}}}(\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}}))\in\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}\text{ for all }b_{{\mathbf{x}}}\in\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}}.$
Proof..
We will make the standard identification of ${\mathcal{T}}_{\lambda}(X)$ with $\overline{\pi}({\mathcal{T}}_{\lambda}(X))$, and of ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ with $\mathrm{C}^{*}(\overline{\pi},U)$. In order to make a distinction we write $e_{G}$ for the unit of $G$ and $e_{\mathfrak{H}}$ for the unit of ${\mathfrak{H}}$. For the first part note that $\iota^{\rtimes}_{\ast}(\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}})=\mathbf{K}_{{\mathbf{x}},\iota^{\rtimes}_{\ast}}$ since $\iota^{\rtimes}$ is a representation of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Moreover by checking the generators we have that
$\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}\subseteq\mathbf{K}_{{\mathbf{x}},\iota^{\rtimes}_{\ast}}.$
For the reverse inclusion, let $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ such that $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e_{G}$ and $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P={\mathbf{x}}$, and consider an element $b_{\mathbf{x}}$ in $\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}}$ of the form:
$(\hat{t}^{\rtimes}\left(\overline{\pi}\left(\lambda_{p_{1}}(\xi_{p_{1}})\right)U_{{\mathfrak{h}}_{1}}\right)^{*})^{\varepsilon}\hat{t}^{\rtimes}(\overline{\pi}\left(\lambda_{q_{1}}(\xi_{q_{1}})\right)U_{{\mathfrak{h}}_{2}})\cdots\hat{t}^{\rtimes}(\overline{\pi}\left(\lambda_{p_{n}}(\xi_{p_{n}})\right)U_{{\mathfrak{h}}_{2n-1}})^{*}\hat{t}^{\rtimes}(\overline{\pi}\left(\lambda_{q_{n}}(\xi_{q_{n}})\right)U_{{\mathfrak{h}}_{2n}})^{\varepsilon^{\prime}}.$
By the covariance relation of $(\overline{\pi},U)$, and the fact that $\alpha_{\mathfrak{h}}(\lambda_{p}(X_{p}))=\lambda_{p}(X_{p})$ for all $p\in P$ and ${\mathfrak{h}}\in{\mathfrak{H}}$, we can write
$\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}})$ $=(\left(\overline{\pi}\left(\lambda_{p_{1}}(\xi_{p_{1}})\right)U_{{\mathfrak{h}}_{1}}\right)^{*})^{\varepsilon}\overline{\pi}\left(\lambda_{q_{1}}(\xi_{q_{1}})\right)U_{{\mathfrak{h}}_{2}}\cdots(\overline{\pi}\left(\lambda_{p_{n}}(\xi_{p_{n}})\right)U_{{\mathfrak{h}}_{2n-1}})^{*}(\overline{\pi}\left(\lambda_{q_{n}}(\xi_{q_{n}})\right)U_{{\mathfrak{h}}_{2n}})^{\varepsilon^{\prime}}$
$=\overline{\pi}\left(\lambda_{p_{1}}(\eta_{p_{1}})^{*}\right)^{\varepsilon}\overline{\pi}\left(\lambda_{q_{1}}(\eta_{q_{1}})\right)\cdots\overline{\pi}\left(\lambda_{p_{n}}(\eta_{p_{n}})^{*}\right)\overline{\pi}\left(\lambda_{q_{n}}(\eta_{q_{n}})\right)^{\varepsilon^{\prime}}U_{{\mathfrak{h}}}$
$=\overline{\pi}\left(\left(\lambda_{p_{1}}(\eta_{p_{1}})^{*}\right)^{\varepsilon}\lambda_{q_{1}}(\eta_{q_{1}})\cdots\lambda_{p_{n}}(\eta_{p_{n}})^{*}\lambda_{q_{n}}(\eta_{q_{n}})^{\varepsilon^{\prime}}\right)U_{{\mathfrak{h}}}$
for ${\mathfrak{h}}:={\mathfrak{h}}_{1}^{-\varepsilon}{\mathfrak{h}}_{2}\cdots{\mathfrak{h}}_{2n-1}^{-1}{\mathfrak{h}}_{2n}^{\varepsilon^{\prime}}$, and some appropriate choice of $\eta_{p_{i}}\in X_{p_{i}}$ and $\eta_{q_{i}}\in X_{q_{i}}$ for $i=1,\dots,n$. Therefore we have that
$\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}})\in\overline{\pi}(\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}})\cdot U_{{\mathfrak{h}}}\subseteq\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
By considering linear combinations and their norm-limits we have that
$\mathbf{K}_{{\mathbf{x}},\iota^{\rtimes}_{\ast}}\subseteq\mathbf{K}_{\mathbf{x},\lambda_{\ast}}\rtimes_{\alpha,\lambda}{\mathfrak{H}},$
as required.
For the second part, let $p_{1},q_{1},\dots,p_{n},q_{n}\in P$ and $\varepsilon,\varepsilon^{\prime}\in\{0,1\}$ such that $p_{1}^{-\varepsilon}q_{1}\cdots p_{n}^{-1}q_{n}^{\varepsilon^{\prime}}=e_{G}$ and $q_{n}^{-\varepsilon^{\prime}}p_{n}\dots q_{1}^{-1}p_{1}^{\varepsilon}P={\mathbf{x}}$, and consider an element $b_{\mathbf{x}}$ in $\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}}$ of the form:
$(\hat{t}^{\rtimes}\left(\overline{\pi}\left(\lambda_{p_{1}}(\xi_{p_{1}})\right)U_{{\mathfrak{h}}_{1}}\right)^{*})^{\varepsilon}\hat{t}^{\rtimes}(\overline{\pi}\left(\lambda_{q_{1}}(\xi_{q_{1}})\right)U_{{\mathfrak{h}}_{2}})\cdots\hat{t}^{\rtimes}(\overline{\pi}\left(\lambda_{p_{n}}(\xi_{p_{n}})\right)U_{{\mathfrak{h}}_{2n-1}})^{*}\hat{t}^{\rtimes}(\overline{\pi}\left(\lambda_{q_{n}}(\xi_{q_{n}})\right)U_{{\mathfrak{h}}_{2n}})^{\varepsilon^{\prime}}.$
Due to the crossed product covariance and the definition of the action we can write
$\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}})=\left(\overline{\pi}\left(\lambda_{p_{1}}(\eta_{p_{1}})\right)^{*}\right)^{\varepsilon}\overline{\pi}\left(\lambda_{q_{1}}(\eta_{q_{1}})\right)\cdots\overline{\pi}\left(\lambda_{p_{n}}(\eta_{p_{n}})\right)^{*}\overline{\pi}\left(\lambda_{q_{n}}(\eta_{q_{n}})\right)^{\varepsilon^{\prime}}U_{{\mathfrak{h}}}$
for ${\mathfrak{h}}:={\mathfrak{h}}_{1}^{-\varepsilon}{\mathfrak{h}}_{2}\cdots{\mathfrak{h}}_{2n-1}^{-1}{\mathfrak{h}}_{2n}^{\varepsilon^{\prime}}$, and some $\eta_{p_{i}}\in X_{p_{i}}$ and $\eta_{q_{i}}\in X_{q_{i}}$ for $i=1,\dots,n$. By definition we have that
$E_{{\mathfrak{H}}}(\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}}))=\begin{cases}\overline{\pi}\left(\left(\lambda_{p_{1}}(\eta_{p_{1}})^{*}\right)^{\varepsilon}\lambda_{q_{1}}(\eta_{q_{1}})\cdots\lambda_{p_{n}}(\eta_{p_{n}})^{*}\lambda_{q_{n}}(\eta_{q_{n}})^{\varepsilon^{\prime}}\right)&\text{if }{\mathfrak{h}}=e_{{\mathfrak{H}}},\\ \hskip 2.84544pt0&\text{otherwise},\end{cases}$
and thus $E_{{\mathfrak{H}}}(\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}}))\in\overline{\pi}\left(\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}\right)$ as required. The proof is completed by considering finite linear combinations of elements of this form in $\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}}$, and their norm-limits. ∎
The following proposition is the key result of this section.
Proposition 4.5.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\alpha$ be a generalised gauge action of a discrete group ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$. Then the identity representation $\iota^{\rtimes}\colon X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\to{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is equivariant, Fock covariant, and injective.
Proof..
We will show that the identity representation satisfies the conditions of Theorem 3.2. By definition it is injective. Moreover, by [50, Lemma 7.16] we have that
$({\mathcal{T}}_{\lambda}(X)\otimes\mathrm{C}^{*}_{\max}(G))\rtimes_{\alpha\otimes{\operatorname{id}},\lambda}{\mathfrak{H}}\simeq({\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}})\otimes\mathrm{C}^{*}_{\max}(G),$
by a canonical *-isomorphism, and hence we get that $\iota^{\rtimes}$ is equivariant. It remains to show that $\iota^{\rtimes}$ satisfies conditions (i) and (ii) of Theorem 3.2. We will make the standard identification of ${\mathcal{T}}_{\lambda}(X)$ with $\overline{\pi}({\mathcal{T}}_{\lambda}(X))$, and of ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ with $\mathrm{C}^{*}(\overline{\pi},U)$.
For condition (i), an application of Proposition 4.4 and using that $\mathbf{K}_{\emptyset,\lambda_{\ast}}=(0)$ yield
$\mathbf{K}_{\emptyset,\iota^{\rtimes}_{\ast}}=\mathbf{K}_{\emptyset,\lambda_{\ast}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}=(0).$
For condition (ii), let =${\mathcal{F}}=\{{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}\}$ be a finite $\cap$-closed subset of ${\mathcal{J}}$ such that $\bigcup_{i=1}^{n}\mathbf{x}_{i}\neq\emptyset$ and let $b_{{\mathbf{x}}_{i}}\in\mathbf{K}_{{\mathbf{x}}_{i},\hat{t}^{\rtimes}_{\ast}}$ for $i=1,\dots,n$ that satisfy
$\sum_{i:r\in{\mathbf{x}}_{i}}\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}_{i}})\iota^{\rtimes}_{r}\left(\overline{\pi}\left(\lambda_{r}(X_{r}\right)U_{{\mathfrak{h}}}\right)=(0)\text{ for all }r\in\bigcup_{i=1}^{n}{\mathbf{x}}_{i}\text{ and}{\mathfrak{h}}\in{\mathfrak{H}}.$
We will show that $\sum_{i=1}^{n}\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}_{i}})=0$. For notational convenience, set
$c:=\sum_{i=1}^{n}\iota^{\rtimes}_{\ast}(b_{{\mathbf{x}}_{i}})\quad\text{and}\quad c_{r}:=\sum_{i:r\in\mathbf{x}_{i}}\iota^{\rtimes}_{\ast}(b_{\mathbf{x}_{i}})\text{ for each }r\in\bigcup_{i=1}^{n}{\mathbf{x}}_{i}.$
Since $b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}}\in\mathbf{K}_{{\mathbf{x}}_{i}\cap{\mathbf{x}}_{j},\hat{t}^{\rtimes}_{\ast}}$, by Proposition 4.4 we have that
$E_{{\mathfrak{H}}}\left(\iota^{\rtimes}_{\ast}(b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}})\right)\in\overline{\pi}(\mathbf{K}_{{\mathbf{x}}_{i}\cap{\mathbf{x}}_{j},\lambda_{\ast}}),$
and in particular we can write
$E_{{\mathfrak{H}}}\left(\iota^{\rtimes}_{\ast}(b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}})\right)=\overline{\pi}\left(\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})\right),\text{ for some }d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}}\in\mathbf{K}_{\mathbf{x}_{i}\cap\mathbf{x}_{j},\hat{t}_{\ast}}.$
Therefore we have that
$E_{{\mathfrak{H}}}(c^{*}c)=\sum_{i,j=1}^{n}E_{{\mathfrak{H}}}(\iota^{\rtimes}_{\ast}(b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}}))=\overline{\pi}\left(\sum_{i,j=1}^{n}\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})\right)\in\overline{\pi}\left(\mathbf{B}_{{\mathcal{F}},\lambda_{\ast}}\right),$
where we used that $\{\mathbf{x}_{i}\cap\mathbf{x}_{j}:i,j=1,\dots,n\}={\mathcal{F}}$ for the inclusion, since ${\mathcal{F}}$ is $\cap$-closed. Now let
$r\in\bigcup_{i,j=1}^{n}(\mathbf{x}_{i}\cap\mathbf{x}_{j})=\bigcup_{i=1}^{n}\mathbf{x}_{i}.$
By using that ${\mathcal{F}}$ is $\cap$-closed we obtain that
$\{(i,j)\in[n]^{2}\mid r\in\mathbf{x}_{i}\cap\mathbf{x}_{j}\}=\{i\in[n]\mid r\in\mathbf{x}_{i}\}\times\{j\in[n]\mid r\in\mathbf{x}_{j}\},$
and hence
$E_{{\mathfrak{H}}}(c_{r}^{*}c_{r})$ $=\sum_{i:r\in\mathbf{x}_{i}}\sum_{j:r\in\mathbf{x}_{j}}E_{{\mathfrak{H}}}\left(\iota^{\rtimes}_{\ast}(b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}})\right)=\sum_{i,j:r\in\mathbf{x}_{i}\cap\mathbf{x}_{j}}E_{{\mathfrak{H}}}\left(\iota^{\rtimes}_{\ast}(b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}})\right)=\sum_{i,j:r\in\mathbf{x}_{i}\cap\mathbf{x}_{j}}\overline{\pi}\left(\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})\right).$
Since $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, by using the assumption we have that $c_{r}\overline{\pi}(\lambda_{r}(\xi_{r}))=0$ for all $\xi_{r}\in X_{r}$. Thus we get
$\overline{\pi}(\lambda_{r}(\xi_{r}))^{*}E_{{\mathfrak{H}}}(c_{r}^{*}c_{r})\overline{\pi}(\lambda_{r}(\xi_{r}))=E_{{\mathfrak{H}}}\left(\overline{\pi}(\lambda_{r}(\xi_{r}))^{*}c_{r}^{*}c_{r}\overline{\pi}(\lambda_{r}(\xi_{r}))\right)=0,$
where we used that ${\mathcal{T}}_{\lambda}(X)$ is in the multiplicative domain of $E_{\mathfrak{H}}$. Consequently, we get that $E_{{\mathfrak{H}}}(c_{r}^{*}c_{r})^{1/2}\overline{\pi}(\lambda_{r}(\xi_{r}))=0$ and in particular
$\overline{\pi}\left(\sum_{i,j:r\in\mathbf{x}_{i}\cap\mathbf{x}_{j}}\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})\lambda_{r}(\xi_{r})\right)=\sum_{i,j:r\in\mathbf{x}_{i}\cap\mathbf{x}_{j}}\overline{\pi}\left(\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})\right)\overline{\pi}(\lambda_{r}(\xi_{r}))=E_{{\mathfrak{H}}}(c_{r}^{*}c_{r})\overline{\pi}(\lambda_{r}(\xi_{r}))=0.$
Since $\overline{\pi}$ is faithful we deduce that
$\sum_{i,j:r\in\mathbf{x}_{i}\cap\mathbf{x}_{j}}\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})\lambda_{r}(X_{r})=(0)\text{ for all }r\in\bigcup_{i,j=1}^{n}(\mathbf{x}_{i}\cap\mathbf{x}_{j}).$
Since $\lambda$ satisfies condition (ii) of Theorem 3.2 we get that $\sum_{i,j=1}^{n}\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})=0$, and therefore
$E_{{\mathfrak{H}}}(c^{*}c)=\overline{\pi}\left(\sum_{i,j=1}^{n}\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})\right)=0.$
Faithfulness of $E_{{\mathfrak{H}}}$ then implies that $c=0$, as required. ∎
Since $A\times_{X,\lambda}P$ is the C*-envelope of ${\mathcal{T}}_{\lambda}(X)^{+}$ it inherits a group action $\dot{\alpha}$ from ${\mathfrak{H}}$. The reduced Hao–Ng isomorphism problem asks if there exists a canonical *-isomorphism such that
$(A\rtimes_{\alpha,\lambda}{\mathfrak{H}})\times_{X\rtimes_{\alpha,\lambda}{\mathfrak{H}},\lambda}P\stackrel{{\scriptstyle?}}{{\simeq}}(A\times_{X,\lambda}P)\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}.$
By [48, Theorem 5.1] this is equivalent to asking for a canonical *-isomorphism
(4.2) $\mathrm{C}^{*}_{\textup{env}}({\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{+})\stackrel{{\scriptstyle?}}{{\simeq}}\mathrm{C}^{*}_{\textup{env}}({\mathcal{T}}_{\lambda}(X)^{+})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}},$
which by Proposition 4.2 is equivalent to the existence of a canonical completely isometric isomorphism
(4.3) ${\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{+}\simeq{\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
A careful investigation of the arguments in [11, 12, 24, 25, 28] suggests that it is enough to show that the embedding $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\hookrightarrow{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is Fock covariant, without passing through a *-isomorphism of their ambient C*-algebras.
Theorem 4.6.
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\alpha$ be a generalised gauge action of a discrete group ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$. Then the identity representation $\iota^{\rtimes}\colon X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\to{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ lifts to a completely isometric isomorphism
${\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{+}\simeq{\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}}.$
Consequently, the reduced Hao–Ng isomorphism problem has an affirmative answer, i.e.,
$(A\rtimes_{\alpha,\lambda}{\mathfrak{H}})\times_{X\rtimes_{\alpha,\lambda}{\mathfrak{H}},\lambda}P\simeq(A\times_{X,\lambda}P)\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}},$
by a canonical *-isomorphism, where $\dot{\alpha}$ is the induced action of ${\mathfrak{H}}$ on $A\times_{X,\lambda}P$.
Proof..
By Proposition 4.5 the identity representation is an equivariant Fock covariant injective representation of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Moreover, by [50, Lemma 7.16] we have that
$({\mathcal{T}}_{\lambda}(X)\otimes\mathrm{C}^{*}_{\lambda}(G))\rtimes_{\alpha\otimes{\operatorname{id}},\lambda}{\mathfrak{H}}\simeq({\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}})\otimes\mathrm{C}^{*}_{\lambda}(G),$
by a canonical *-isomorphism, and hence we get that the coaction that $\iota^{\rtimes}$ admits is normal. Since
${\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}}=\overline{\operatorname{alg}}\{\iota^{\rtimes}_{p}(X_{p}\rtimes_{\alpha,\lambda}{\mathfrak{H}})\mid p\in P\},$
the required completely isometric isomorphism is induced by Corollary 4.1. By Proposition 4.2 we then have that
$\mathrm{C}^{*}_{\textup{env}}({\mathcal{T}}_{\lambda}(X\rtimes_{\alpha,\lambda}{\mathfrak{H}})^{+})\simeq\mathrm{C}^{*}_{\textup{env}}({\mathcal{T}}_{\lambda}(X)^{+})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}},$
and [48, Theorem 5.1] finishes the proof. ∎